### v3

# Beyond Any-Shot Adaptation: Predicting Optimization Outcome for Robustness Gains without Extra Pay

Abstract
The foundation model enables general-purpose problem-solving and enjoys desirable rapid adaptation due to its adopted cross-task generalization paradigms, e.g., pretraining, meta-training, and finetuning. Recent advances in these paradigms show the crucial role of challenging tasks’ prioritized sampling in enhancing adaptation robustness. However, ranking task difficulties exhausts massive task queries to evaluate, thus computation and annotation intensive, which is typically unaffordable in practice. This work underscores the criticality of both adaptation robustness and learning efficiency, especially in scenarios where tasks are risky or costly to evaluate, e.g., policy evaluations in Markov decision processes (MDPs) or inference with large models. To this end, we present Model Predictive Task Sampling (MPTS) to establish connections between the task space and adaptation risk landscape to form a theoretical guideline in robust active task sampling. MPTS characterizes the task episodic information with a generative model and directly predicts task-specific adaptation risk values from posterior inference. The developed risk learner can amortize expensive evaluation and provably approximately rank task difficulties in the pursuit of task robust adaptation. MPTS can be seamlessly integrated into zero-shot, few-shot, and many-shot learning paradigms. Extensive experimental results are conducted to exhibit the superiority of the proposed framework, remarkably increasing task adaptation robustness and retaining learning efficiency in contrast to existing state-of-the-art (SOTA) methods. The code is available at the project site https://github.com/thu-rllab/MPTS.
1 Introduction
Figure 1: Framework of Model Predictive Task Sampling in Any-Shot Learning. a. The task episodic learning might involve MDPs, large models, or annotators in task construction for optimization. The adaptive machine learner can be a large model or RL policy. MPTS contains posterior inference, adaptation outcome simulation, and task subset selection for any-shot learning. [Snow: freeze models; Fire: update models] b. MPTS amortizes the process, e.g., task batch construction, environment interaction, or large models’ inference at a lower computational budget with the risk learner. c. The risk learner takes in the coupled identifier and the exact adaptation risk to train under streaming variational inference. d. The risk learner simulates the adaptation outcomes of $\hat{\mathcal{B}}$ candidate identifiers, computes acquisition scores, and selects $-\mathcal{B}$ ones for next iteration.
Enhancing machine learning models’ generalization across various scenarios is a long-standing pursuit in artificial general intelligence. The rise of generative artificial intelligence provides a promising solution to this problem, which facilitates the development of foundation models 1, 2, 3. Unlike traditional models, which are specific to learning tasks and cannot sufficiently meet the demands of time-sensitive responses, e.g., online intelligent recommendation or consultancy, the foundation model exhibits the advantages of fast deployment in a wide range of tasks without learning from scratch.
One secret behind foundation models’ success lies in rapid adaptation acquired from pretraining, meta-training, and finetuning. Take meta-learning as an example; it optimizes the machine learner over the task distribution and shares a similar cognition in human’s developing skills 4, where we consolidate cumulated experience into knowledge and quickly convert the few-shot guidelines to a specific skill. Technically, adaptation learning paradigms update machine learners episodically and leverage extracted problem-solving prior to unseen but related scenarios in a zero-shot, few-shot, or many-shot manner 1, 5. Such a mechanism exploits slow learning on massive tasks to solve new tasks fast at test time.
Research Motivations:
Currently, there have been several studies in enhancing adaptation, e.g., the neural scaling law in foundation models 6, 7 implies that cross-task generalization scales with model complexity, task capacity, and computational power. Recognized as a few-shot learner, large language models (LLMs) mandate the collection of large datasets as tasks to train for generalization 8. In robotics, a popular schema for generalist policies is diversifying the decision-making environments, e.g., Markov decision processes (MDPs), for the robot to interact and execute policy optimization 9. As a consequence, task episodic learning is commonly used in learning to adapt over task distributions, where the task batch is randomly resampled, e.g., from a uniform random sampler, to optimize over iteration (see Fig. 1/2b). Meanwhile, these distributions are induced by the task identifiers; e.g., in Fig. 2a, diverse physics parameters are used to create diverse MDPs in domain randomization (DR) 9 and meta reinforcement learning (Meta-RL) 10.
Meanwhile, as distribution shifts 11 exist ubiquitously at the task level in the real world, adaptation robustness is catching increasing attention in the field 12, 13. Some trivial yet frequently adopted sampling strategies, e.g., random sampling from a uniform task distribution, are prone to encounter catastrophic failure in adaptation to risk-sensitive scenarios due to undersampling critical tasks. For example, traffic accidents are more informative than usual cases when training autonomous driving systems but occupy a small portion 14. This inspires some advances in challenging tasks’ prioritized sampling strategies 14, 15, 16, which necessitates task difficulty ranking during optimization. For the sake of improving adaptation robustness, these approaches 17, 18, 16, 19, 20 rank the task samples after evaluation and prioritize difficult ones to optimize (see evaluate-rank-filter in Table 1). Nevertheless, exactly evaluating tasks from either losses or gradients brings extra computational and annotation overhead. In LLMs, the hallucination stands as a lasting challenge, and the post-training phase sometimes relies on prohibitively intensive manual evaluation and precise human feedback such as millions of interactions, which serves for effective alignment 21, 22. In DR or Meta-RL, the agent has to interact with numerous MDPs to collect after-adaptation episodes and compute returns. All of these raise concerns about learning efficiency during optimization for robust adaptation, particularly when the machine learner employs a large backbone or interactions with environments are expensive.
Encouraged by the pressing real-world necessities, we delve into robust active task sampling, which can potentially cut off extra costs arising from task construction, intensive annotations, or computations when evaluating the machine learner’s adaptation to tasks. In the presence of any-shot adaptation, it is desirable to seek a task sampling strategy that requires fewer learning resources but retains more deployment benefits such as robust pattern recognition with foundation models and risk-averse sequential decision-making.
Developed Approach:
Our brain can simulate the outcome of some unencountered scenarios based on continually cumulated experience without exactly putting actions into practice. The inherent cognitive system is also robust to handle extreme cases and energy efficient when developing diverse skills through an implicit task selection mechanism. Hence, it prompts us to learn to model the optimization process of any-shot adaptation and simultaneously adjust the task sampling strategy from the predicted outcome as feedback. A similar rationale is reflected in neuroscience 4, 23, 24; the prefrontal cortex resembles a meta-reinforcement learning system, implying the adaptation feedback also interleaves with the task query process. As a result, this work investigates the design of task robust sampling strategies with the guidance of a risk predictive model. We posit key insights in methodology: (i) Some variables, i.e., adaptation risk, are predictable during episodic learning, which is informative to rank task difficulties and steer task selection; (ii) Generative modeling task-specific adaptation risk enables capturing risk landscapes with quantified uncertainty, which aligns optimization processes with certain robustness principles.
To this end, we develop a new framework oriented to task episodic learning, namely Model Predictive Task Sampling (MPTS) to enable robust active task sampling in risk minimization. As displayed in Fig. 1a, our strategy exploits the historical risk information to construct a lightweight surrogate model referred to as the risk learner, foresighting adaptation risk across the task space for the task sampler and guiding the optimization of the adaptive machine learner. Specifically, we view actively selecting task batches as planning from past experience with the assistance of the risk learner in MPTS. Fig. 1b reveals the risk learner’s role in amortizing the expensive evaluation process on an arbitrary set of tasks conditioned on an adaptive machine learner, which further serves task difficulty scoring. The risk learner in Fig. 1c is in a variational autoencoder (VAE) structure 25 and engages in online adaptation risk generation via posterior inference. The acquisition function in Fig. 1d comprises worst-case performance and the predictive uncertainty into selection criteria to rate the task identifier candidate for filtering purposes.
MPTS also draws inspiration from active inference 26, which involves perception, action, and learning phases in a loop for uncertainty minimization about the planning environment. Similarly, this work conceptually characterizes task episodic learning by a risk generative model in Fig. 2b-c, and the subset selection from the task batch can be treated as online planning to derive a robust machine learner. Technically, MPTS specifies or infers the identifiers from the task distribution (see some instantiations in Fig. 2a) to establish mappings inside the risk learner and adopts the streaming variational inference (VI) in training the model. In addition, running acquisition functions in a larger identifier batch to pick up the task subset for the next iteration balances exploration (uncertainty minimization) and exploitation (worse-case robustness) in the task space. In primary, our proposed MPTS enjoys several benefits in robust optimization:
  1. 1.
    Adaptation Robustness. The optimization pipeline of MPTS can advance the machine learner’s adaptation robustness in the presence of severe task distribution shifts, such as subpopulation shifts and domain generalization;
  2. 2.
    Learning Efficiency. As the lightweight risk learner amortizes expensive evaluation parts, MPTS can diminish computational overhead, avoid unnecessary annotations, and encourage exploration in the task space;
  3. 3.
    Framework Versatility. Leveraging the optimization episodes, MPTS can work as a plug-play module to rank the task difficulties in optimization and allow seamless integration into robust zero-shot, few-shot, and many-shot learning.
This work performs extensive experiments in diverse task episodic learning scenarios, including few-shot regression, few-shot image classification with foundation models, Metra-RL, robotic DR, and prompt-tuning foundation models. Evidence indicates that incorporating MPTS into any-shot adaptation learning can significantly promote adaptation robustness in most scenarios. Compared with existing SOTA robust adaptation methods, MPTS raises no extra computation, memory usage, or interactions with environments and even accelerates the learning process under certain scenarios.
Figure 2: Fundamental Concepts: Task Identifiers, Episodic Learning and Probabilistic Graphical Models. a. The task distribution is uniform and defined over meaningful identifiers $\bm{\tau}$. For example, the amplitude and the phase $[a,b]$ specifies a sinusoid curve to complete with K-shot observed data points. Robots like Half-Cheetahs are trained to accomplish different locomotion tasks with varying masses and velocities. Some multimodal pattern recognition tasks’ identifiers are implicit but can be described from a reference model, e.g., text encoders in CLIP 1. b. The episodic learning in the task space resamples a batch of tasks from $p(\tau)$ to optimize in each iteration. c. Here, the generative model includes grey units as observed variables and white ones as unobservable. The solid directed lines describe the generative model 27. We use the dash-directed lines to indicate the recognition model and approximate inference within autoencoding variational Bayes 25.
2 Adaptation and Robustness
Notations.
We denote a task sample by $\tau\sim p(\tau)$, with $\mathcal{T}$ the task space. The identifier, a real-value vector $\bm{\tau}$ illustrated in Fig. 2a, specifies the task $\tau$ in the distribution. Let the task-specific risk function be $\ell:\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}\times\Theta\mapsto\mathbb{R}$, which evaluates $\bm{\theta}$ parameterized machine learner’s adaptation performance on the task $\tau$. Take few-shot regression as an instance, $\mathcal{D}_{\tau}^{S}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=1}^{K}$ refers to the support dataset available for fast adaptation, e.g., finetuning to obtain the model $p_{\bm{\theta}}(\bm{y}|\mathcal{D}_{\tau}^{S},\bm{x})$; while $\mathcal{D}_{\tau}^{Q}=\{[\bm{x}_{i},\bm{y}_{i}]\}_{i=K+1}^{K+N}$ denotes the query dataset to evaluate the after-adaptation machine learner’s performance, e.g., $\ell=-\frac{1}{N}\sum_{i=1}^{N}\ln p_{\bm{\theta}}(\bm{y}_{i}|\mathcal{D}_{\tau}^{S},\bm{x}_{i})$.
Given the condition $|\mathcal{D}_{\tau}^{S}|=\emptyset$, the risk function $\ell$ examines the zero-shot adaptation capability of a trained model. Otherwise, $\ell$ reflects the few-shot or many-shot adaptation risk (e.g., supervised finetuning over many examples). During optimization, we can write the history of the episodic task batch information as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$ with $\mathcal{B}$ the number of tasks in one batch, where $\bm{\theta}_{t}$ represents the machine learner’s parameter in the $t$-th iteration.
The tuple set $\left\{\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$ involves the sampled task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$, the query and support dataset $\{\mathcal{D}_{\tau_{t,i}}:=\mathcal{D}_{\tau_{t,i}}^{S}\cup\mathcal{D}_{\tau_{t,i}}^{Q}\}_{i=1}^{\mathcal{B}}$, and the evaluated adaptation risk $\{\ell_{t,i}\}_{i=1}^{\mathcal{B}}$ in the $t$-th iteration. For convenience, we express the simplified risk history as $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$, which depends on $\bm{\theta}_{t}$.
2.1 Adaptation Risk Function and Task Filtering
The task episodic learning setup optimizes the machine learning within $p(\tau)$. Our analysis is interested in the risk landscape in the task space as illustrated in Fig. 1d. Such a perspective emphasizes the interplay between the task identifier $\bm{\tau}$, the task-specific dataset $\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$ and the adaptation risk function $\ell$ conditioned on $\bm{\theta}$. To account for varying levels of adaptation, we delineate three distinct scenarios: zero-shot, few-shot, and many-shot learning. As finetuning is treated as a many-shot adaptation, we skip the instantiation below.
Zero-Shot Adaptation Risk Function.
During training, we evaluate $\ell$ on the query dataset $\mathcal{D}_{\tau}^{Q}$ conditioned on the machine learner $\bm{\theta}$, i.e., $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$. With robotic DR 9 as an example, $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta})$ denotes the negative return of trajectories collected under the policy $\bm{\theta}$. This evaluation is commonly seen in zero-shot learning without support information.
Few-shot Adaptation Risk Function.
The form of $\ell$ is specific to meta-learning methods. For instance, MAML 28 implements a bi-level optimization framework. In this context, $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})$ is written as $\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta}_{\text{meta}}-\alpha\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{S};\bm{\theta}))$, where $\bm{\theta}_{\text{meta}}$ denotes the meta initialization, and the inside-bracket gradient term corresponds to finetuning the model tailored to $\tau$ with $\alpha$ the learning rate.
Task Filtering in Robust Adaptation Learning Pipelines.
Recent works 29, 19, 30 uncover compelling evidence that filtering some tasks to train under several criteria, e.g., ‘easy,’ ‘difficult,’ or ‘learnable’ in evaluation can enhance machine learners’ adaptation robustness and even improve task efficiency compared with uniform sampling strategies. The ‘difficult’ prioritization criteria aligns with minimizing conditional value-at-risk ($\text{CVaR}_{\alpha}$) 31, a metric in evaluating robustness 32, 16, 33. These methods operate under an Evaluate-Rank-Filter mode for the next model update in Table 1 and induce the batch filtering ratio $\hat{\alpha}=1-\frac{\mathcal{B}}{\hat{\mathcal{B}}}\in[0,1)$, which describes the fraction of discarded tasks in a sample batch.
Table 1: A Summary of Typical Risk Minimization Approaches. The previously mentioned ERM 34, GDRM 35, 36, 37, DRM 16, 33, 29, 30, and MPTS are included in comparison. In DRM, the filtering criteria is to discard proportional easy tasks in a batch.
        Principle         Robustness Concept         Optimization Pipeline
        ERM         No         Evaluate $\Rightarrow$ Update
        GDRM         Uncertainty Set         Evaluate $\Rightarrow$ Reweight $\Rightarrow$ Update
        DRM         Tail Task Risk         Evaluate $\Rightarrow$ Rank $\Rightarrow$ Filter $\Rightarrow$ Update
        MPTS (Ours)         Tail Task Risk with Uncertainty         Simulate $\Rightarrow$ Rank $\Rightarrow$ Filter $\Rightarrow$ Update
2.2 Risk Minimization Principles
The risk minimization principles are entangled with task sampling and robust optimization.
Expected/Empirical Risk Minimization (ERM).
With the fixed $p(\tau)$, the principle follows the statistical learning theory 34 and minimizes the expectation of adaptation risk over the task space. As a result, we can have:
$\min_{\bm{\theta}\in\bm{\Theta}}\mathbb{E}_{p(\tau)}\Big{[}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\Big{]}.$ (1)
It draws batches with a random task sampler to optimize iteratively.
Group Distributionally Robust Risk Minimization (GDRM).
Such a principle 35 effectively improves robustness in distribution shifts and has shown positive effects on training foundation models 36, 37. It constructs a collection of uncertainty sets over tasks and results in the optimization objective as follows:
$\begin{split}\min_{\bm{\theta}\in\bm{\Theta}}\sup_{g\in\mathcal{G}}\mathbb{E}_{p_{g}(\tau)}\Big{[}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\Big{]},\end{split}$ (2)
where $\mathcal{G}$ are groups of uncertainty sets, and $p_{g}(\tau)$ indicates the probability measure over the task group. The operation inside Eq. (2) prioritizes the worst group to optimize in a soft way.
Distributionally Robust Risk Minimization (DRM).
We retain the notation of task robust works 16, 38, which terms the tail task risk minimization as DRM. It aims to improve the robustness of adaptation to the tail tasks over iteration. No explicit form exists as the tail task distribution is $\bm{\theta}$-dependent. The optimization objective is derived as the $\text{CVaR}_{\alpha}(\bm{\theta})$ 31:
$\min_{\bm{\theta}\in\bm{\Theta}}\text{CVaR}_{\alpha}(\bm{\theta}):=\mathbb{E}_{p_{\alpha}(\tau;\bm{\theta})}\Big{[}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\Big{]},$ (3)
where we write $p_{\alpha}(\tau;\bm{\theta})$ to express the $(1-\alpha)$ proportional worst case for easier formulation. Also note that when $\alpha$ approaches $1$, the problem degenerates to the worst-case risk minimization.
3 Results
This section reports theoretical and empirical findings in robust adaptation and analyzes the effect of MPTS in task episodic learning. Prior to elaborating on the experimental setups, we outline the workflows underpinning MPTS.
Optimization Outcome Prediction with Theoretical Guarantee & MPTS Guided Risk Minimization.
Given the task episodic learning setup and notations, we simply describe the optimization process in any-shot adaptation:
[…]
Prior robust adaptation works 14, 15, 16, 17, 20 require evaluating the machine learner over tasks for further robust optimization while the task batch size $\mathcal{B}$ is restricted by the sample or memory cost. Meanwhile, nearly all of them fail to utilize the above optimization outcomes $H_{1:t}$; our finding is that under Assumptions 1/2/3, the cumulated risk episodes can be used to train a risk predictive model, i.e., the risk learner in MPTS, through posterior inference and work as the informative prior for continually scoring each task’s adaptation difficulty.
Assumption 1 (Lipschitz Continuity)
We assume the adaptation risk function $\ell(\cdot;\bm{\theta})$ reserves the Lipschitz continuity w.r.t. $\bm{\theta}$ and $\bm{\tau}$, i.e.,
$\begin{split}|\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})-\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}^{\prime})|\leq\beta_{1}||\bm{\theta}-\bm{\theta}^{\prime}||\quad\text{and}\quad|\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})-\ell(\mathcal{D}_{\tau^{\prime}}^{Q},\mathcal{D}_{\tau^{\prime}}^{S};\bm{\theta})|\leq\beta_{2}||\bm{\tau}-\bm{\tau}^{\prime}||,\end{split}$ (5)
where $\forall\{\bm{\theta},\bm{\theta}^{\prime}\}\in\bm{\Theta}$ and $\forall\{\bm{\tau},\bm{\tau}^{\prime}\}\in\mathcal{T}$ with Lipschitz constants $\beta_{1}$ and $\beta_{2}$.
Assumption 2 (Bounded Sample Gradient)
We assume the norm of the adaptation risk function’s gradient $\nabla\ell(\cdot;\bm{\theta}_{t})$ is bounded:
$\sup_{\tau\in\mathcal{T}}||\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}_{t})||_{2}<G_{t},$ (6)
where $G_{t}$ is a positive constant.
Assumption 3 (Sub-Gaussian Stochastic Gradient)
The stochastic gradient $\tilde{\bm{g}}:=\bm{g}+\bm{\epsilon}$ for the machine learner’s adaptation at $t$-th iteration is $\sigma$-sub-Gaussian, which means:
$\begin{split}\mathbb{E}\left[\exp{(\eta\bm{v}^{T}\bm{\epsilon})}\right]\leq\exp{\left(\frac{\eta^{2}\sigma^{2}||\bm{v}||_{2}^{2}}{2}\right)}\quad\forall\eta\in\mathbb{R}\ \text{and}\ \bm{v}\in\mathbb{R}^{d},\end{split}$ (7)
where $\mathbb{E}[\tilde{\bm{g}}]=\bm{g}$, $\mathbb{E}[||\tilde{\bm{g}}-\bm{g}||_{2}^{2}]\leq\sigma^{2}$ and $\sigma\in\mathbb{R}^{+}$.
Theorem 1 (Provably Approximately Invariant Task Difficulties)
Given arbitrary $K$ data points $\{(\bm{\tau}_{i},\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})\}_{i=1}^{K}$, the adaptation gradient $\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})$ as a $\sigma$-sub-Gaussian random variable and $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})$, we denote the relative difficulty via the difference $\Delta_{ij}(\bm{\theta}_{t+1})=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t+1})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t+1})$ and $\Delta_{ij}(\bm{\theta}_{t})=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})$ between $t$-th and $(t+1)$-th iterations, and the gradient difference as $\bm{v}_{ij}:=\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})$.
Under Assumption 1/2/3, the set of rank-preserving variable $E_{ij}:=\mathds{1}\left[\text{sign}(\Delta_{ij}(\bm{\theta}_{t+1}))=\text{sign}(\Delta_{ij}(\bm{\theta}_{t}))\right]$ satisfies the probability inequality:
$\mathbb{P}(\bigcap_{i<j}E_{ij})\geq 1-\xi,$
when $\eta\leq\frac{\delta_{t}}{2G_{t}M_{t}+\sqrt{8\sigma^{2}G_{t}^{2}\ln\left(\frac{K(K-1)}{2\xi}\right)}}$ with $G_{t}$ in Assumption 2, $\delta_{t}:=\min_{i\neq j}|\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})|\in\mathbb{R}^{+}$, the stochastic gradient norm $M_{t}:=||\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})||_{2}$.
Let us first track adaptation risk episodes as the approximate prior for MPTS to online learn. Coupling the identifier $\bm{\tau}$ and adaptation risk $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})$ in the task batch constructs the risk history $H$, which serves as the streaming database for training the risk learner. The continuity assumption of $\ell(\cdot;\bm{\theta})$ w.r.t. the identifier $\bm{\tau}$ makes it reasonable to predict the optimization outcome of an arbitrary task based on risk histories, i.e., $p(\ell|\bm{\tau},H_{1:t};\bm{\theta}_{t})$ $\forall\bm{\tau}\in\mathcal{T}$. Meanwhile, learning stochastic adaptation risk function $p(\ell|\bm{\tau},H_{1:t};\bm{\theta}_{t})$ provides a probabilistic landscape of risk statistics over iterations.
MPTS amortizes arbitrary task adaptation performance evaluation and approximately ranks task difficulties. Learning to predict the optimization outcome $p(\ell|\bm{\tau},H_{1:t};\bm{\theta}_{t})$ is hopeful for performance evaluation in infinite tasks with negligible computations and enlarges the pseudo batch size $\hat{\mathcal{B}}$ for filtering, implicitly encouraging exploration in the task space. Theorem 1 provides a provable foundation for optimization outcome prediction, which suggests that a controllable perturbation over $\bm{\theta}$, such as a one-step gradient update with a smaller enough learning rate, changes the overall difficulties between tasks with a minor probability. Hence, the candidate tasks $\mathcal{T}_{t+1}^{\hat{\mathcal{B}}}$ evaluated by the latest machine learner $\bm{\theta}_{t}$ in posterior inference probably preserves their relative task difficulties given the updated $\bm{\theta}_{t+1}$. The following empirical results will confirm such predictability and its difficulty scoring ability with MPTS.
For ease of presentation, this work treats MPTS as a variant of the risk minimization principle under specific acquisition criteria. As illustrated in Fig. 1 and Fig. 2b-c, the core workflow of MPTS contains continually training a risk predictive model $p(\ell|\bm{\tau},H_{1:t};\bm{\theta}_{t})$, evaluating task-specific adaptation risk through posterior inference, and then screening the task batch under the upper confidence bound (UCB) principle 39 for $(t+1)$-th optimization. Accordingly, these operations constitute key steps in Eq. (8), where Monte Carlo estimate by the risk learner is used to obtain the mean $m(\ell)$ and the standard deviation $\sigma(\ell)$ of arbitrary task’s adaptation risk and the acquisition function $\mathcal{A}(\cdot)$ quantifies the uncertainty-aware average performance in the task subset.
[…]
Approximating optimization outcome after adaptation relies on streaming VI 40, 41, and the risk learner is a lightweight model. By leveraging the historical risk information, MPTS surrogates partial pipelines of DRM and amortizes the computationally heavy, annotation-intensive, or task-query expensive components. This design not only enhances learning efficiency but also aligns with the overarching goals of risk-averse adaptation. Hence, MPTS can be viewed as a surrogate for task distributionally robust methods, such as $\text{CVaR}_{1-\mathcal{B}/\hat{\mathcal{B}}}$ minimization; repeating the boxed steps until convergence leads to the robust adaptive machine learner. The details of MPTS and implementations are attached in Methods part (Section 5).
Benchmark of Any-shot Adaptation.
The experimental design takes the benchmark typicality and the practical challenges into consideration. Testing scenarios span downstream tasks across pattern recognition and decision-making to assess MPTS, with certain experiments involving multimodal foundation models. These experiments mainly examine few-shot adaptation and include (1) K-shot sinusoid regression 28, (2) N-way K-shot image classification 42 with CLIP models and (3) Meta-RL 28. Additionally, MPTS explores the application scenarios like (4) robotic DR 43, 9 for zero-shot adaptation and (5) CLIP model’s parameter-efficient finetuning (PEFT) towards image classification.
Backbones and Task Robust Baselines.
Note that the primary focus of this study lies in boosting adaptation robustness; the comparison baselines are in accordance with the previously mentioned risk minimization principles. While these principles, along with our MPTS, are agnostic to specific zero-shot, few-shot, or many-shot learning methods, we adopt the representative or SOTA method as the default backbone in the experiment.
In sinusoid regression and Meta-RL, we employ MAML 28 as the backbone algorithm. In N-way K-shot image classification, CLIP 1 has already achieved remarkable zero-shot performance, and we utilize its prompt-based extension MaPLe 44 for classification. In robotic DR for zero-shot adaptation, we set TD3 45 as the base policy optimization algorithm due to its stable performance. In many-shot adaptation, we still use MaPLe 44 for prompt-tuning in image classification.
To implement baselines, we incorporate Table 1 listed ERM 34, DRM 16, 33, GDRM 35 and MPTS (ours) into the above backbone methods. In terms of adaptation robustness, we compute $\text{CVaR}_{\alpha}$ across validation task set or testing task datasets with $\alpha=\{0.9,0.7,0.5\}$, with some OOD results reported. Moreover, we evaluate these methods from efficiency metrics, such as computational cost, memory usage, and sample efficiency in separate scenarios. For a fair comparison, the task batch to optimize $\mathcal{B}$ is the same for all baselines, excluding those easier tasks pruned after evaluation. Unless notable clarifications, the sampled task batch size for ERM and GDRM is $\mathcal{B}$ while that for DRM is ^$\hat{\mathcal{B}}=2\mathcal{B}$ with half filtered for stable optimization. We refer the reader to Supplementary Notes Section F/G for configuration details.
Figure 3: K-shot Sinusoid Regression Results (7 Runs). a. Shown are curves of averaged MSEs on the validation task set during meta-training for all methods . b. The meta-trained machine learners are tested on a fixed task set, reporting the average MSEs and CVaR values. c. Displayed are meta-testing results with MPTS machine learners trained by various $\gamma_{1}/\ \gamma_{0}$ ratios. d. Displayed are meta-testing results with MPTS machine learners trained in various pseudo batch sizes, i.e., ^=$\hat{\mathcal{B}}=\{1\mathcal{B},2\mathcal{B},4\mathcal{B},8\mathcal{B}\}$. e. The PCC values are tracked during meta-training. f. At a specific iteration, the statistical correlation between predicted and exact adaptation risk values of the task batch is visualized with overall $\rho_{\bar{\ell},\ell}=0.669$. g. The required relative run-time is computed for all methods during meta-training with ERM as the anchor. h. At some meta-training time-step, we visualize the process of next-batch selection from the pseudo batch and the risk learner.
3.1 Demonstration of the MPTS’s role in K-shot sinusoid regression
In K-shot sinusoid regression, the function family $\{f(x)=a_{i}\sin(x-b_{i})|(a_{i},b_{i})\in[0.1,5.0]\times[0.0,\pi]\}$ is specified by the identifier $\bm{\tau}=[a,b]$. We take this as a toy demonstration to gain insight into MPTS and the role of the risk learner.
MPTS allows for roughly scoring the task difficulty with risk learners over iterations.
In challenging tasks’ prioritized sampling 14, 15, 16, 33, assessing task difficulties is identified as the preliminary step. Note that the risk learner learns to amortize the exact evaluation $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}_{t})$ $\forall\bm{\tau}\in\mathcal{T}\ \text{and}\ \bm{\theta}_{t}\in\bm{\Theta}$ with the help of risk histories, and the evaluation result implicitly reflects adaptation difficulties. Once the risk learner can roughly judge difficulties between tasks, we can trust the amortized result in MPTS for the follow-up robust optimization.
Hence, given $(t+1)$-th screened task batch and its evaluation, we keep track of the risk learner’s predicted adaptation risk values, {¯$\{\bar{\ell}_{t+1,i}:\approx\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}[p_{\bm{\psi}}(\ell|\bm{\tau}_{t+1,i},H_{1:t})]\}_{i=1}^{\mathcal{B}}$ and the corresponding exact ones $\{\ell_{t+1,i}\}_{i=1}^{\mathcal{B}}$ from $\bm{\theta}_{t+1}$ for computing the Pearson correlation coefficient (PCC) value $\rho_{\bar{\ell},\ell}:=\frac{\sum_{i=1}^{\mathcal{B}}(\bar{\ell}_{t+1,i}-\text{Mean}[\{\bar{\ell}_{t+1,.}\}])(\ell_{t+1,i}-\text{Mean}[\{\ell_{t+1,.}\}])}{\sqrt{\sum_{i=1}^{\mathcal{B}}(\bar{\ell}_{t+1,i}-\text{Mean}[\{\bar{\ell}_{t+1,.}\}])^{2}}\sqrt{\sum_{i=1}^{\mathcal{B}}(\ell_{t+1,i}-\text{Mean}[\{\ell_{t+1,.}\}])^{2}}}$ over iterations. The predicted risk values are actually amortized evaluation results conditioned on $\bm{\theta}_{t}$, but it will not affect empirical conclusion from PCC values due to the provably relative difficulty relationship preservation in Theorem 1. Under this consideration, PCC values reasonably quantify the task difficulty ranking capability of the risk learner in a rough granularity.
As shown in Fig. 3e, $\rho_{\bar{\ell},\ell}$ consistently ranges between 0.4 and 0.8 across iterations, validating the reliability of the risk learner’s stochastic forward passes for forecasting adaptation outcomes in batches. We observe a gradual decline in $\rho_{\bar{\ell},\ell}$ over iterations, a pattern also detected across other experiments. This decline likely stems from the convergence of the machine learner $\bm{\theta}_{t}$ and task difficulty stabilization within $\mathcal{T}$ during later training stages, where diminished diversity between successive task batches negatively impacts the risk learner’s training dynamics. Fig. 3f illustrates the predicted and exact adaptation batch risks and their statistical correlations at a specific iteration. The scattered points demonstrate strong overall alignment between predicted and exact risk values, even though the value scale between iterations varies a lot. Notably, difficult tasks with persistently high MSEs along the $x$-axis cluster around or above the correlation slope.
The risk learner accelerates the learning process and improves comphrehensive adaptation performance with the help of amortized evaluation.
Fig. 3a shows that MPTS achieves quick convergence, completing optimization after 15K iterations compared to 20K required by ERM and GDRM. This acceleration benefits from the uncertainty-guided worst-case acquisition strategy. DRM processes $\mathcal{B}$ tasks per iteration and selects half in optimization, bringing a 1.7x computational overhead relative to ERM (see Fig. 3g). In contrast, the involvement of the risk learner in MPTS incurs a 0.14x runtime increase, a marginal over ERM. To intuitively understand active task sampling in MPTS, we visualize the predicted adaptation risk values spanned over the task space and the selected task subset from the candidate identifiers in Fig. 3h. The selected task identifiers from the pseudo batch favor regions with higher deviations, and the majority of them concentrate on high-risk regions.
In meta-testing, Fig. 3b confirms the MPTS and DRM’s lowest average and $\text{CVaR}_{\alpha}$ mean-square errors (MSEs), and their advantages over GDRM and ERM grow significant with increasing confidence level $\alpha$. As empirically verified in work 46, DRM typically sacrifices task and computational efficiency to enhance the robustness of machine learners, often requiring intensive task evaluation to filter. For the used MAML backbone, the gradient update as the inner loop for adaptation further increases the computational overhead, while MPTS implicitly bypasses it through probabilistic prediction.
The appropriate hyper-parameter configuration secures performance and efficiency.
We first examine the acquisition function $\mathcal{A}(\mathcal{T}^{\mathcal{B}};\bm{\phi},\bm{\psi})$’s influence through the trade-off parameters $\{\gamma_{0},\gamma_{1}\}$ in Eq. (16). To this end, we meta-test machine learners’ average performance trained with various setups $\frac{\gamma_{1}}{\gamma_{0}}=\{1.0,3.0,6.0,9.0\}$, $\gamma_{0}=0.0$ and $\gamma_{1}=0.0$ in Fig. 3c. The results depict that relatively higher uncertainty weights can decrease the average MSEs to a certain level while simply eliminating the worst-case consideration with $\gamma_{0}=0.0$ weakens the performance.
Moreover, we scrutinize the influence of the pseudo batch size $\hat{\mathcal{B}}$ in Fig. 3d, where testing models are trained with diverse $\hat{\mathcal{B}}$. Increasing $\hat{\mathcal{B}}$ to a certain scale significantly reduces the average meta-testing MSEs in Fig. 3d. Nevertheless, excessively larger $\hat{\mathcal{B}}$, e.g., ^$\hat{\mathcal{B}}=8\times\mathcal{B}$, adversely affects testing performance. This is because the selected task identifiers might concentrate under the worst-case selection principle when increasing the number of identifiers and over-optimizing the local region of the task space deteriorates the global generalization.
Overall, the appropriate configuration of MPTS adheres to the following disciplines: (i) $\hat{\mathcal{B}}$ greater than $\mathcal{B}$ should be set within an acceptable size to encourage task exploration and prevent excessive optimization within a local task region; (ii) since adaptation robustness improvement remains the primary goal, we consistently set $\gamma_{0}\in\mathbb{R}^{+}$ as the default across all experiments.
Figure 4: 5-way 1-shot Meta-testing Classification Results (3 Runs in Average). a-d. Shown are testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with meta-trained machine learners on different datasets. e. With experiments on ImageNet-A as an example, we report the memory cost and clock time relative to ERM during meta-training for all methods.
3.2 Few-Shot adaptation benefits from MPTS in robustness and learning efficiency
Result analysis in N-way K-shot image classification.
We perform 5-way 1-shot image classification using MaPLe. The six meta-training datasets are respectively constructed by ImageNet-CG 47, ImageNet-CI 47, ImageNet-CS 47, ImageNet-A 48, ImageNet-S 49 and ImageNet-R 50 in a standard manner. Explicit $\bm{\tau}$ are unavailable to specify the task; however, it can be approximately resolved by describing the identifier through a small reference model. Specifically, we leverage CLIP’s text encoder to obtain $\bm{\tau}\approx[\text{CLIP}_{\text{text}}(\mathcal{C}_{1}),\dots,\text{CLIP}_{\text{text}}(\mathcal{C}_{K})]$ with the tokenizations of $K$ class texts $\mathcal{C}_{1:K}$ from $\mathcal{D}_{\tau}^{S}$.
During meta-training, resource consumption on large image datasets is quantified in Fig. 4e, comparing relative computational time and memory usage across methods. The extra memory and computation cost arising from optimizing risk learners can be negligible in MPTS compared with the total ones. DRM demands relatively 1.3x computational time and 1.6x memory consumption compared with ERM, as $2\mathcal{B}$ tasks are fed into the model to evaluate and filter half in each iteration.
In meta-testing, MPTS achieves the highest average accuracies on six image datasets in Fig. 4d and Table 5. As for robustness, the results in Fig. 4a-c witness a comprehensively increase of $\text{CVaR}_{0.5}$, $\text{CVaR}_{0.7}$ and $\text{CVaR}_{0.9}$ values in both MPTS and DRM. Each specific method’s performance exhibits alike patterns across different datasets. All robust optimization methods show robustness improvement over ERM to some extent. Among them, MPTS and DRM are leading in all metrics, while the latter requires more learning resources to guarantee adaptation robustness. Overall, difficult task prioritization in this benchmark can simultaneously increase average performance as a side product.
Figure 5: Meta-RL Results on Five Mujoco Environments (7 Runs). a. The cumulative returns with standard error of means (SEMs) belonging to $\text{CVaR}_{0.9}$ validation MDPs are displayed during meta-training. b. We compute the average cumulative returns with SEMs on validation MDPs during meta-training. c. Tracked are the risk learner’s PCC values with SEMs over training iterations. d. The relative clock time quantifies the computational complexity for all methods on Walker2dVel, where ERM’s runtime works as the anchor. e. We report $\text{CVaR}_{\alpha}$ returns of meta-testing MDPs. f. The box-plot reports results averaged over meta-testing MDPs.
Result analysis in Meta-RL.
We first analyze the meta-training results. As displayed in Fig. 2a, the identifier refers to specific physics variables, explicit for risk learners’ construction. In Fig. 5a, on all benchmarks except ReacherPos, MPTS reaches the highest $\text{CVaR}_{0.9}$ validation returns, followed by DRM and GDRM. ERM largely underperforms MPTS, which is consistent on all benchmarks. In terms of computational and sample efficiency, DRM takes about 1.5x computational time on Walker2dVel due to extra environment interactions under inner gradient updated policies to filter tasks, while MPTS does not suffer this issue (see Fig. 5d). In Fig. 5b, DRM sacrifices more average returns than ERM on HalfCheetahMassVel, HalfCheetahVel, and Walker2dVel. Though MPTS can be viewed as the surrogate for DRM, the involvement of active sampling and extra task exploration makes MPTS’s average performance comparable with ERM on HalfCheetahMassVel and Walker2dVel. GDRM behaves mediate throughout evaluations in Fig. 5a-b. DRM mostly strikes a balance between the average and $\text{CVaR}_{0.9}$ performance; however, DRM and MPTS receive both the average and $\text{CVaR}_{0.9}$ bonus on ReacherPos. This reflects that the effect of worst task selection might vary between environments. The tracked $\rho_{\bar{\ell},\ell}$ in Fig. 5c reveals the risk learner’s strong capability of discriminating MDP difficulties, staying above 0.5 nearly across all environments over iterations. The trend of $\rho_{\bar{\ell},\ell}$ on ReacherPos differs from those on other environments and does not decline along iteration.
Meta-testing results in Fig. 5e-f displays performance differences across CVaR values and other baselines’ robustness improvements over ERM, which align with the learning curves in Fig. 5a. MPTS’s performance advantage notably amplifies with increasing $\alpha$ values. In the extreme worst case $\text{CVaR}_{0.9}$, MPTS beats ERM by more than 20% returns on all benchmarks. Average meta-testing performance across environments draws no consistent conclusions. Walker2dMassVel and Walker2dVel show minor differences between methods (see Fig. 5f), while on HalfCheetahMassVel, MPTS achieves comparable returns to ERM despite slightly higher variance. HalfCheetah verifies the marginal advantages of GDRM and ERM, whereas ReacherPos favors MPTS and DRM with reduced variance. Importantly, MPTS, serving as a DRM proxy, essentially maintains average performance comparable with ERM.
In total, MPTS enjoys significant adaptation robustness benefits across most scenarios with mediate average performance, but it reserves more computational and sample efficiency than DRM. Unlike few-shot classification cases, Meta-RL mostly trades off the worst-case and average performance, similarly observed in work 33.
3.3 MPTS retains multi-faced advantages beyond robustness in zero-shot continuous control
DR is a popular paradigm for enabling zero-shot skill transfer in robotics 9. Most setups in DR resemble meta-reinforcement learning, except that no support episodes are available for few-shot adaptation.
Figure 6: DR Results on Ergo-Reacher and Lunar-Lander (7 Runs). a. In Ergo-Reacher, the $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average cumulative returns on validation MDPs are reported together with the risk learner’s PCC curve during DR training. b. In Lunar-Lander, the cumulative returns on validation MDPs are illustrated together with the risk learner’s PCC curve during DR training. c. We test the DR-trained policies on the fixed MDP set and report the $\text{CVaR}_{\alpha}$ cumulative returns. d. The returns averaged over DR-testing MDPs are illustrated. e. The required runtime is computed for all methods on Lunar-Lander. f. In Lunar-Lander, shown are frequencies of sampled identifiers using MPTS during DR training. g. In Lunar-Lander, we test the trained policies in both in-distribution (ID) domains and out-of-distribution (OOD) domains to report each task’s average returns.
MPTS dominates the overall performance in DR training.
The Ergo-Reacher curves oppose general Meta-RL findings; the use of robust optimization, such as MPTS and DRM, improves both the average and CVaR performance in Fig. 6a. One probable reason is that MPTS benefits from extra task exploration with the relatively larger $\hat{\mathcal{B}}$. Meanwhile, $\rho_{\bar{\ell},\ell}$ fluctuates near 0.4 throughout training (see Fig. 6a). In Lunar-Lander, Fig. 6b shows that MPTS can keep the leading trend in both average and CVaR performance. However, DRM and GDRM not only fall behind others in the average returns but also receive the lowest $\text{CVaR}_{0.9}$ returns, completely failing in robust optimization. The simple worst-case task selection or reweighted mechanism might deteriorate overall performance when some challenging tasks unavailable to be successfully resolved are yet sampled with a high frequency. Instead, MPTS enlarges the search scope and selects a subset based on the rule hybrid of the worst-case and the uncertainty, which avoids over-optimizing the finite worst MDPs. Regarding $\rho_{\bar{\ell},\ell}$, it quickly peaks above 0.6 and stabilizes around 0.3 over iterations. This is in accordance with prior evidence; the selected tasks around the convergence iteration might concentrate and weaken the discrimination of task difficulties.
In Lunar-Lander, the runtime of MPTS is comparable with ERM and GDRM (see Fig. 6e). Moreover, we record the frequency of sampling different MDPs during DR training, and the histogram in Fig. 6f indicates that MPTS reserves more preference for MDPs with smaller engine strength while still exploring MDPs with all possible main engine strengths at the early stage.
MPTS facilitates zero-shot policy adaptation in the worst-case and OOD MDPs.
Given the DR-trained policies, we first test CVaR and average returns in ID tasks in Fig. 6c-d. In Ergo-Reacher, MPTS and DRM achieve the highest CVaR, and ERM surprisingly obtains a faint average performance’ dip than others in the boxplot. In Lunar-Lander, MPTS retains the highest CVaR returns, and its returns even hardly decrease with $\alpha$ up. MPTS reaches a notable surge over ERM by more than 20%. Besides, MPTS and ERM achieve top-average returns with the smallest deviations in the boxplot.
In Lunar-Lander, $\bm{\tau}\in\mathbb{R}^{+}$ means the main engine strength of the rocket. We shift the DR-training identifier’s interval from $\bm{\tau}\in[4.0,20.0]$ to $\bm{\tau}\in[1.0,4.0)\cup(20.0,23.0]$ and test DR-trained policies on these OOD MDPs. Compared with the tested ID MDP returns (see the middle of Fig. 6g), all methods encounter sharp performance decline in the hard OOD tasks and retain faint return fluctuations in the easy OOD tasks $\bm{\tau}\in(20.0,23.0]$. DRM-trained policies exhibit more deviations and fail to advance generalization to easier tasks. However, MPTS still dominates returns over others in the hard OOD tasks, exhibiting strong adaptation in difficult cases.
3.4 MPTS also reserves the potential of robust PEFT
Figure 7: Testing Classification Results after Prompt-Tuning on ImageNet (3 Runs in Average). a-d. Shown are testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with the prompt-tuned machine learner on ID and OOD datasets. e. During prompt-tuning ImageNet, we report the memory cost and clock time relative to ERM for all methods.
In terms of finetuning scenarios, we can view each labeled example in the dataset as a task. A general implementation can be found in Algorithm 6/7. Here, prompt tuning, as a recent PEFT method, is used to learn the prompt token for adapting the pretrained model with many-shot examples in testing scenarios. We retain all implementation in MaPLe 44, execute PEFT on the ImageNet 51 tuning dataset, and then perform standard evaluation. To examine the prompt-tuned model’s robustness on the OOD dataset, we conduct domain generalization and test it on four other datasets with diverse domain shifts, including ImageNet-A 48, ImageNet-S 49, ImageNet-R 50, and ImageNet-V 52.
In Fig. 7a-d, MPTS consistently outperforms other baselines in average and different CVaR accuracies on ID and OOD testing datasets. MPTS beats ERM by 0.82%-3.11% accuracy increase in terms of $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$ and $\text{CVaR}_{0.5}$ indicators (see Table 6). MPTS is more advantageous on OOD datasets than ImageNet results. On 4/5 datasets except ImageNet-A, DRM is nearly SOTA secondary to MPTS on CVaR accuracies but reaches ERM comparable performance in average accuracies. GDRM’s performance varies with $\alpha$ levels and shows marginal gains compared with ERM. Like the few-shot image classification case, DRM again sacrifices the computational and memory efficiency for specific robustness improvement as reported in Fig. 7e. Even though MPTS shares a similar optimization purpose with DRM, the risk predictive module and simulation over larger batches empower MPTS with more task exploration power at negligible computational cost, resulting in a more optimal robust machine learner.
4 Discussion
Rapid adaptation to novel scenarios with any-shot examples has been identified as an indispensable capability in artificial general intelligence. While the neural scaling laws stress the importance of the scale of task datasets, practical considerations, such as safety issues and limited annotation and computational budgets, necessitate a closer focus on adaptation robustness and learning efficiency. This study takes a closer look at the optimization process of any-shot adaptation through generative modeling and presents MPTS as a versatile module for robust active task sampling.
Primary Findings.
First of all, this work reports a vital finding that it is theoretically feasible to predict the optimization outcome in a rough granularity for the online scoring difficulty of any task for the machine learner to adapt. Extensive experimental results demonstrate that MPTS can (i) amortize any-shot adaptation evaluation for robustness purposes and promote exploration in the task space, (ii) significantly enhance adaptation robustness across diverse tasks, including pattern recognition and decision-making, and (iii) accelerate the learning process under scenarios like few-shot regression and DR.
Future Extensions.
This work’s main contribution is a novel framework of task sampling to promote robustness without sacrificing learning efficiency. The large-scale experimental results show the promise of risk predictive models in surrogating adaptation performance evaluation, which can be integrated with more task selection rules in developing the foundation models. However, more effort is needed to scale MPTS to training or tuning foundation models with more complicated multimodal signals.
5 Methods
In alignment with the realistic necessities, this work focuses on robust adaptation while securing learning efficiency, such as circumventing expensive evaluation to allow for real-world adaptation. Such a purpose facilitates the birth of MPTS. As previously mentioned, the framework is agnostic to task episodic learning methods; hence, we leave out any-shot learning details. As noticed in Fig. 1a, several roles are involved in the optimization: (1) the adaptive machine learner, e.g., foundation models or generalist policies, learns to learn in an any-shot way given some optimizers; (2) the risk learner as a critic evaluates and forecasts the task-specific adaptation risk; (3) the task sampler as an actor works for screening task batches for next iteration. These components participate in episodic learning until convergence.
Technically, this work recasts task episodic learning to sequence generation and develops MPTS as the task sampling strategy to balance exploration and exploitation. To reconcile theory and practice, we first introduce a tractable optimization approach to enable functional posterior inference towards adaptation risk. Then, we devise the acquisition function informed by the captured risk landscapes. Finally, interpretations concerning the optimization pipeline are attached to conclude the method part.
5.1 Generative Modeling Risk Functions and Posterior Inference
Here, we design the sampling strategy through the lens of risk landscapes and pay more attention to datasets of learning optimization outcome $\{H_{t}\}_{t=1}^{T}$. To characterize the adaptation risk during batch optimization, we introduce the latent variable $\bm{z}_{t}$ to summarize episodic information and present a versatile deep generative model as:
$\begin{split}p(H_{0:T},\bm{z}_{0:T}|\bm{\theta}_{0:T})=p(\bm{z}_{0})\prod_{t=0}^{T}p_{\bm{\psi}}(H_{t}|\bm{z}_{t};\bm{\theta}_{t})\prod_{t=0}^{T-1}p(\bm{z}_{t+1}|\bm{z}_{t}).\end{split}$ (9)
Within a Bayesian framework, we approximate the underlying function distribution with the latent variable, and the posterior $p(\bm{z}_{t}|H_{t})$ summarizes the historical risk information and accounts for uncertainty in distributions. The following writes the form of $p(\bm{z}_{t}|H_{t})$ according to the Bayes rule 53:
[…]
where $p(\bm{z}_{t}|H_{1:t-1})$ encodes the past evaluation results as the conditional prior. Moreover, $p(H_{t}|\bm{z}_{t})$ conveys the likelihood of producing observations of the task batch risk values in the $t$-th iteration. Notably, the exact computation w.r.t. the posterior is intractable due to the complicated integral in the denominator.
Generative Process.
As illustrated in Fig. 2c, risk values of the task batch $\ell$ are correlated with the meta learner’s parameters $\bm{\theta}$. In specific, the factorization of the sequential optimization relevant variables arrives at:
[…]
where $\bm{z}_{t}$ in the probabilistic graphical model constitutes the distribution over risk functions (For the sake of simplicity, we skip over other variables less relevant to our learning purposes). Here, we assume the conditional independence between task-specific risk values given $\bm{z}$ and the machine learner’s parameter $\bm{\theta}$ in Eq. (11). And the primary optimization objective is to $\max_{\bm{\psi}\in\bm{\Psi}}\ln p_{\bm{\psi}}(H_{t}|H_{1:t-1})$ for the optimization outcome prediction.
Inference Process.
The manner of episodic training, where the task batch and its evaluation arrive sequentially, inspires us to predict adaptation risk values online to actively sample tasks in a batch. However, the exact inference w.r.t. $p(\bm{z}_{t}|H_{t})$ is infeasible as there is no structural information regarding posteriors. In each iteration, the risk function distribution relies on the updated machine learner $\bm{\theta}$; hence, such non-stationarity in the risk function distributions prompts us to involve the streaming VI 40, 41 to derive the approximate posterior.
To do so, we handle the streaming task batches and update the posterior in a recursive way:
$\begin{split}\underbrace{p(\bm{z}_{t}|H_{t})}_{\text{{Updated Posterior}}}\propto\underbrace{p(H_{t}|\bm{z}_{t})}_{\text{{Likelihood}}}\underbrace{p(\bm{z}_{t}|H_{1:t-1})}_{\text{{Functional Prior}}}\end{split}$ (12)
where $p(\bm{z}_{t}|H_{1:t-1})$ represents the conditional prior using the last time updated posterior as the proxy. The role of the estimated functional posterior is to provide uncertainty-aware prediction and serves the task sampling strategy design, which will be detailed in Section 5.2.
As a result, we can formulate the evidence lower bound (ELBO) as a tractable optimization objective in Eq. (13) from approximate inference.
$\begin{split}\small\max_{\bm{\psi}\in\bm{\Psi},\bm{\phi}\in\bm{\Phi}}\hat{\mathcal{G}}_{\text{ELBO}}(\bm{\psi},\bm{\phi}):=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\left[\sum_{i=1}^{\mathcal{B}}\ln p_{\bm{\psi}}(\ell_{t,i}|\bm{\tau}_{t,i},\bm{z}_{t})\right]-D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel p(\bm{z}_{t}|H_{1:t-1})\Big{]}\end{split}$ (13)
For implementation convenience, we adopt the parameterized Gaussian distribution with diagonal covariance matrices as variational distributions similar to vanilla VAEs 25, 54 and neural processes (NPs) 55. In other words, these distribution parameters are approximated with neural networks, e.g., $q_{\bm{\phi}}(\bm{z}_{t}|H_{t})=\mathcal{N}(\bm{z}_{t};\mu_{\bm{\phi}}(H_{t}),\Sigma_{\bm{\phi}}(H_{t}))$, and the reparameterization trick 25 is used for stochastic gradient estimate.
As for the neural architecture, we employ the DeepSet encoding module 56 to process the set dataset $H_{t}$, which corresponds to the permutation invariant function family in Definition 3. Also, in the context of streaming VI, $q_{\bm{\phi}}(\bm{z}_{t}|H_{t-1})$ mostly works as the proxy for the conditional prior as default. Consequently, we can modify the exact ELBO in Eq. (13) and further translate the practical optimization process with the Lagrange multiplier $\beta$ as:
$\max_{\bm{\psi}\in\bm{\Psi},\bm{\phi}\in\bm{\Phi}}\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\left[\sum_{i=1}^{\mathcal{B}}\ln p_{\bm{\psi}}(\ell_{t,i}|\bm{\tau}_{t,i},\bm{z}_{t})\right]\quad\text{s.t.}\quad D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel q_{\bar{\bm{\phi}}}(\bm{z}_{t}|H_{t-1})\Big{]}\leq\epsilon\Leftrightarrow$ (14a)
$\max_{\bm{\psi}\in\bm{\Psi},\bm{\phi}\in\bm{\Phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi}):=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\left[\sum_{i=1}^{\mathcal{B}}\ln p_{\bm{\psi}}(\ell_{t,i}|\bm{\tau}_{t,i},\bm{z}_{t})\right]-\beta D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel q_{\bar{\bm{\phi}}}(\bm{z}_{t}|H_{t-1})\Big{]},$ (14b)
where $\bar{\bm{\phi}}$ indicates no gradients computed through $\bm{\phi}$ in the term, and $\{\beta\in\mathbb{R}^{+},\epsilon\in\mathbb{R}^{+}\}$ constrains the machine learner’s parameter search in next iteration.
5.2 Task Sampling Strategy Design
In robust active task sampling, existing strategies evaluate task batches to rank their difficulties in adaptation and then prioritize challenging subsets for optimization 20, 16. Besides the expensive evaluation cost, these strategies are weak in the efficient exploration of the task space.
As Theorem 1 has established the theoretical foundation of amortized evaluation, this necessitates the development of the risk learner from cumulated risk histories. With the model predictive results as amortized evaluation, specific rules can be incorporated into the acquisition function for active sampling.
Meanwhile, it is fascinating for the risk learner to evaluate the machine learner’s adaptation to arbitrarily many tasks with minimal computational cost. Hence, we can easily enlarge the pseudo batch size $\hat{\mathcal{B}}$ for more selection candidates and exploit the epistemic uncertainty from the risk learner, encouraging more exploration in the task space.
Evaluating adaptation performance through stochastic forward passes.
The risk learner and estimated functional posteriors in Eq. (11)/(13) work as tools for the active selection of the task batch. Specifically, the predictive distribution can be depicted as:
[…]
The above predictive distribution $p_{\bm{\psi}}(\ell|\bm{\tau},H_{1:t})$ benefits from the Bayesian modeling and provides a tractable way to roughly assess difficulties of tasks throughout the whole task space.
Rank-Flitering the next task batch to episodically train.
After obtaining $p_{\bm{\psi}}(\ell|\bm{\tau},H_{1:t})$, we draw up a batch sampling strategy on the basis of its quantified statistics. The criteria resembles the acquisition function in classical Bayesian optimization (BO), which includes a collection of available evaluation principles, such as expected improvement 57, output information theoretical index 58 or UCB 59.
However, it is also necessary to clarify that the search space is on the sequential task batch instead of machine learners’ parameters, which differs from the ultimate purpose in BO. Central to our approach is the principle of optimism in the face of uncertainty (OFU) 60. We consider the difficult task’s prioritization for robustness and the epistemic uncertainty as pivotal elements in developing acquisition functions. The grounds behind this idea are that (i) the subset with the worst performance deserves extra attention in optimization for adaptation robustness, and (ii) task regions with high predictive uncertainty tend to be underexplored in the last few iterations.
As a result, we present the acquisition function built on the UCB principle 39:
$\begin{split}\mathcal{A}(\mathcal{T}^{\mathcal{B}};\bm{\phi},\bm{\psi})=\sum_{i=1}^{\mathcal{B}}a(\bm{\tau}_{i})=\sum_{i=1}^{\mathcal{B}}\gamma_{0}\overbrace{m(\ell_{i})}^{\text{{Risk Mean}}}+\gamma_{1}\overbrace{\sigma(\ell_{i})}^{\text{{Epistemic Uncertainty}}},\ \text{where}\ \bm{\tau}_{i}\sim p(\tau)\\ \text{with}\ m(\ell_{i})=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\Big{[}p_{\bm{\psi}}(\ell|\bm{\tau}_{i},\bm{z}_{t})\Big{]}\ \text{and}\ \sigma(\ell_{i})=\mathbb{V}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}^{\frac{1}{2}}\Big{[}p_{\bm{\psi}}(\ell|\bm{\tau}_{i},\bm{z}_{t})\Big{]},\end{split}$ (16)
where $m(\ell_{i})$ and $\sigma(\ell_{i})$ are, respectively, the adaptation risk mean and standard deviations, which can be estimated from multiple stochastic forward passes $\bm{z}_{t}\sim p(\bm{z}_{t}|H_{1:t})$ and $\ell\sim p_{\bm{\psi}}(\ell|\bm{\tau}_{i},\bm{z}_{t})$ using the risk generative model. And $\{\gamma_{0},\gamma_{1}\}$ are hyperparameters to balance considerations.
Then, the Simulate-Rank-Filter operation in Eq. (8)c arrives at the task batch for $(t+1)$-th iteration, i.e., $\mathcal{T}_{t+1}^{\mathcal{B}}=\arg\max_{\mathcal{T}^{\mathcal{B}}\subseteq\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}:|\mathcal{T}^{\mathcal{B}}|=\mathcal{B}}\mathcal{A}(\mathcal{T}^{\mathcal{B}};\bm{\phi},\bm{\psi})$. This characterizes the step of the active subset selection from $\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}$, the randomly sampled identifier candidate set with $|\mathcal{T}^{\hat{\mathcal{B}}}_{t+1}|=\hat{\mathcal{B}}$.
In an implementation, we still perform random sampling from $p(\tau)$ and forecast the task-wise acquisition score $a(\cdot)$ from the risk learner. Candidates in $-\mathcal{B}$ acquisition scores are screened to formulate the task batch $\mathcal{T}_{t+1}^{\mathcal{B}}$ for episodic optimization, as illustrated in Fig. 1d. These steps approximately solve Eq. (8)c and obtain $\mathcal{T}_{t+1}^{\mathcal{B}}$ in a heuristic way.
5.3 Sequentially Optimize the Adaptive Machine Learner
Given the screened $\mathcal{T}_{t+1}$, we execute optimization to update the machine learner’s parameters. The task-specific adaptation risk in $(t+1)$-th iteration is written as $\ell_{t+1,i}(\bm{\theta})$ for the selected task $\bm{\tau}_{i}$. The developed MPTS is agnostic to any-shot learning methods, and the following includes the standard update rule for zero-shot, few-shot, and many-shot scenarios.
Machine Learner Updates in Zero-Shot Adaptation:
The zero-shot setup does not require the support dataset to identify the task. Hence, taking the vanilla DR 43 as an instantiation, we can obtain the update rule as:
$\begin{split}\bm{\theta}_{t+1}=\bm{\theta}_{t}-\frac{\lambda}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{t+1,i}}^{Q};\bm{\theta}_{t}),\end{split}$ (17)
where $\bm{\theta}$ denotes the zero-shot learning model parameter with $\lambda$ the learning rate.
Machine Learner Updates in Few-Shot Adaptation:
Still, we take the typical optimization-based method MAML 28 as an instantiation and provide the update rule as follows:
$\ell_{t+1,i}(\bm{\theta})=\ell(\mathcal{D}_{\tau_{t+1,i}}^{Q};\bm{\theta}^{\text{meta}}_{t}-\lambda_{1,1}\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{t+1,i}}^{S}))$ (18a)
$\bm{\theta}^{\text{meta}}_{t+1}=\bm{\theta}^{\text{meta}}_{t}-\frac{\lambda_{1,2}}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\nabla_{\bm{\theta}}\ell_{t+1,i}(\bm{\theta}),\quad\forall i\in\{1,\dots,\mathcal{B}\}$ (18b)
where $\bm{\theta}^{\text{meta}}$ denotes the meta initialization, and $\lambda_{1,1}$ and $\lambda_{1,2}$ are, respectively, learning rates in the inner and outer loops.
Machine Learner Updates in Many-Shot Adaptation:
Here, we take finetuning pretrained models to downstream tasks 61 as an instantiation. In this case, each data point $[\bm{x},\bm{y}]$ can be viewed as a task with either its embedding $\bm{\tau}$ or $\bm{x}$ as the task identifier. Then the model update rule can be:
$\begin{split}\bm{\theta}_{t+1}=\bm{\theta}_{t}-\frac{\lambda}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\nabla_{\bm{\theta}}\ell([\bm{x}_{t+1,i},\bm{y}_{t+1,i}];\bm{\theta}_{t}),\end{split}$ (19)
where $\{[\bm{x}_{t+1,i},\bm{y}_{t+1,i}]\}_{i=1}^{\mathcal{B}}$ denote the sampled task batch for the $(t+1)$-the iteration.
5.4 Overall Algorithm and Interpretation
For the sake of generality, we write the general form of MPTS in Algorithm 1.
Implementation pipelines in any-shot adaptation.
Putting all the ingredients and optimization recipes together, we can provide some examples of implementation. Still, we take some typical adaptation methods and describe pipelines in the zero-shot, few-shot, and many-shot scenarios. This results in Algorithm 2-7. Since the first iteration in Algorithm 2/4/6 does not involve active sampling, due to no latest history, and the task batch follows the standard random sampling setup.
Input : Task distribution $p(\tau)$; Task batch size $\mathcal{B}$; Candidate batch size $\hat{\mathcal{B}}$; Latest updated $\{\bm{\psi},\bm{\phi}\}$; Latest history $H_{t-1}$; Iteration number $K$; Learning rate $\lambda_{2}$.
Output : Selected task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
// Posterior Inference via Stochastic Gradient Variational Bayes
for $i=1$ to $K$ do
       Perform gradient updates given $H_{t-1}$:
       $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
       $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
end for
// Simulating Zero-shot, Few-shot, and Many-shot Adaptation Results
Randomly sample {^$\{\bm{\hat{}}{\bm{\tau}}_{t,i}\}_{i=1}^{\hat{\mathcal{B}}}$ from $p(\tau)$;
Run amortized evaluation on candidate tasks $\{\delta_{i}:=\gamma_{0}m(\ell_{i})+\gamma_{1}\sigma(\ell_{i})\}_{i=1}^{\hat{\mathcal{B}}}$ in Eq. (16);
Rank $\{\delta_{i}\}_{i=1}^{\hat{\mathcal{B}}}$ and screen $-\mathcal{B}$ values;
Return the screened task batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
Algorithm 1 Model Predictive Task Sampling
Connection with sequential decision-making and control.
Intuitively, MPTS resembles model predictive control (MPC) 62 when treating task sampling under some criteria as an optimal planning problem. In this case, the episodic learning process specifies an underlying dynamical system for MPTS to predict with only one future time step in the simulation to assess the influence of selecting the task batch, and the feedback as exact adaptation risk information further helps improve the episodic risk prediction system. In addition, through the lens of sequential decision-making, we can interpret the optimization pipeline of MPTS from the actor-critic framework in RL 63. In detail, the risk learner works as the critic that predicts adaptation performance in the task $\tau$ given a fixed machine learner. Accordingly, the actor plays the role of selecting the task batch from the acquisition function and then executing the machine learner’s optimization. These two roles are entangled in the MPTS pipeline to achieve robust yet efficient adaptation.
Acknowledgments and Disclosure of Funding
This work is funded by National Natural Science Foundation of China (NSFC) projects with Numbers # 62306326 and # 62495091.
Supplementary Materials for Model Predictive Task Sampling in Any-Shot Adaptation
Input : Task distribution $p(\tau)$; Task batch size $\mathcal{B}$; Learning rate $\lambda_{1}$.
Output : Adapted machine learner $\bm{\theta}$.
Set the initial iteration number $t=1$;
Randomly initialize machine learner $\bm{\theta}$;
Randomly initialize risk learner $\{\bm{\psi},\bm{\phi}\}$;
while not converged do
       Execute Algorithm 3 to access the batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$ and induced $\{\mathcal{D}_{\tau_{t,i}}^{Q}\}_{i=1}^{\mathcal{B}}$;
      // Eval Adaptation Performance
       Compute the task specific adaptation risk $\{\ell_{t,i}:=\ell(\mathcal{D}_{\tau_{t,i}}^{Q};\bm{\theta}_{t})\}_{i=1}^{\mathcal{B}}$;
      Return $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$ as the Input to Algorithm 3;
      // Update Machine Learner
       Perform batch gradient updates:
      $\bm{\theta}_{t+1}\leftarrow\bm{\theta}_{t}-\frac{\lambda_{1}}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\nabla_{\bm{\theta}}\ell_{t,i}$;
      Update the iteration number: $t\leftarrow t+1$;
end while
Algorithm 2 MPTS for DR (Zero-Shot Scenarios)
Input : Task distribution $p(\tau)$; Task batch size $\mathcal{B}$; Candidate batch size $\hat{\mathcal{B}}$; Latest updated $\{\bm{\psi},\bm{\phi}\}$; Latest history $H_{t-1}$; Iteration number $K$; Learning rate $\lambda_{2}$.
Output : Task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
// Functional Posterior Inference
for $i=1$ to $K$ do
       Perform gradient updates given $H_{t-1}$:
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
end for
// Simulating Adaptation Results
Randomly sample {^$\{\bm{\hat{}}{\bm{\tau}}_{t,i}\}_{i=1}^{\hat{\mathcal{B}}}$ from $p(\tau)$;
Run amortized evaluation on candidate tasks $\{\delta_{i}:=\gamma_{0}m(\ell_{i})+\gamma_{1}\sigma(\ell_{i})\}_{i=1}^{\hat{\mathcal{B}}}$ in Eq. (16);
Rank $\{\delta_{i}\}_{i=1}^{\hat{\mathcal{B}}}$ and screen $-\mathcal{B}$ values;
Return the screened task batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
Algorithm 3 Model Predictive Task Sampling
Input : Task distribution $p(\tau)$; Task batch size $\mathcal{B}$; Learning rates: $\{\lambda_{1,1},\lambda_{1,2}\}$.
Output : Meta-trained initialization $\bm{\theta}^{\text{meta}}$.
Set the initial iteration number $t=1$;
Randomly initialize meta learner $\bm{\theta}^{\text{meta}}$;
Randomly initialize risk learner $\{\bm{\psi},\bm{\phi}\}$;
while not converged do
       Execute Algorithm 5 to access the batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$ and $\{\mathcal{D}_{\tau_{t,i}}^{S}\cup\mathcal{D}_{\tau_{t,i}}^{Q}\}_{i=1}^{\mathcal{B}}$;
      // Inner Loop to Fast Adapt
       for $i=1$ to $K$ do
             Compute the task-specific gradient: $\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{t,i}}^{S};\bm{\theta})$;
            Perform gradient updates as fast adaptation:
            $\bm{\theta}_{t}^{i}\leftarrow\bm{\theta}_{t}^{\text{meta}}-\lambda_{1,1}\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{S};\bm{\theta})$;
       end for
      // Outer Loop to Meta-train
       Evaluate fast adaptation performance $\{\ell_{t,i}:=\ell(\mathcal{D}_{\tau_{t,i}}^{Q};\bm{\theta}_{t}^{i})\}_{i=1}^{\mathcal{B}}$;
      Return $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$ as the Input to Algorithm 5;
      Perform meta initialization updates:
      $\bm{\theta}_{t+1}^{\text{meta}}\leftarrow\bm{\theta}_{t}^{\text{meta}}-\frac{\lambda_{1,2}}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\nabla_{\bm{\theta}}\ell_{t,i}$;
      Update the iteration number: $t\leftarrow t+1$;
end while
Algorithm 4 MPTS for Model Agnostic Meta Learning (Few-Shot Scenarios)
Input : Task distribution $p(\tau)$; Task batch size $\mathcal{B}$; Candidate batch size $\hat{\mathcal{B}}$; Latest updated $\{\bm{\psi},\bm{\phi}\}$; Latest history $H_{t-1}$; Iteration number $K$; Learning rate $\lambda_{2}$.
Output : Task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
// Functional Posterior Inference
for $i=1$ to $K$ do
       Perform gradient updates given $H_{t-1}$:
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
end for
// Simulating Adaptation Results
Randomly sample {^$\{\bm{\hat{}}{\bm{\tau}}_{t,i}\}_{i=1}^{\hat{\mathcal{B}}}$ from $p(\tau)$;
Run amortized evaluation on candidate tasks $\{\delta_{i}:=\gamma_{0}m(\ell_{i})+\gamma_{1}\sigma(\ell_{i})\}_{i=1}^{\hat{\mathcal{B}}}$ in Eq. (16);
Rank $\{\delta_{i}\}_{i=1}^{\hat{\mathcal{B}}}$ and screen $-\mathcal{B}$ values;
Return the screened task batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
Algorithm 5 Model Predictive Task Sampling
Input : Task distribution $p(\bm{x})$; Task batch size $\mathcal{B}$; Learning rate $\lambda_{1}$.
Output : Fine-tuned machine learner $\bm{\theta}$.
Set the initial iteration number $t=1$;
Randomly initialize machine learner $\bm{\theta}$;
Randomly initialize risk learner $\{\bm{\psi},\bm{\phi}\}$;
while not converged do
       Execute Algorithm 7 to access the batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$ and $\{[\bm{x}_{t,i},\bm{y}_{t,i}]\}_{i=1}^{\mathcal{B}}$;
      // Eval Adaptation Performance
       Compute the instance-specific adaptation risk $\{\ell_{t,i}:=\ell([\bm{x}_{t,i},\bm{y}_{t,i}];\bm{\theta}_{t})\}_{i=1}^{\mathcal{B}}$;
      Return $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$ as the Input to Algorithm 7;
      // Update Machine Learner
       Perform batch gradient updates:
      $\bm{\theta}_{t+1}\leftarrow\bm{\theta}_{t}-\frac{\lambda_{1}}{\mathcal{B}}\sum_{i=1}^{\mathcal{B}}\nabla_{\bm{\theta}}\ell_{t,i}$;
      Update the iteration number: $t\leftarrow t+1$;
end while
Algorithm 6 MPTS for Pretrained Model Finetuning (Many-Shot Scenarios)
Input : Offline processed $\bm{\tau}$ dataset; Task batch size $\mathcal{B}$; Candidate batch size $\hat{\mathcal{B}}$; Latest updated $\{\bm{\psi},\bm{\phi}\}$; Latest history $H_{t-1}$; Iteration number $K$; Learning rate $\lambda_{2}$.
Output : Task identifier batch $\{\bm{\tau}_{t,i}\}_{i=1}^{\mathcal{B}}$.
// Functional Posterior Inference
for $i=1$ to $K$ do
       Perform gradient updates given $H_{t-1}$:
      $\bm{\phi}\leftarrow\bm{\phi}+\lambda_{2}\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
      $\bm{\psi}\leftarrow\bm{\psi}+\lambda_{2}\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ in Eq. (28);
end for
// Simulating Adaptation Results
Randomly sample {^$\{\bm{\hat{}}{\bm{\tau}}_{t,i}\}_{i=1}^{\hat{\mathcal{B}}}$ from $p(\tau)$;
Run amortized evaluation on candidate tasks $\{\delta_{i}:=\gamma_{0}m(\ell_{i})+\gamma_{1}\sigma(\ell_{i})\}_{i=1}^{\hat{\mathcal{B}}}$ in Eq. (16);
Rank $\{\delta_{i}\}_{i=1}^{\hat{\mathcal{B}}}$ and screen $-\mathcal{B}$ values;
// Exact Evaluation or Active Annotations
Return the screened batch $\{[\bm{x}_{t,i},\bm{y}_{t,i}]\}_{i=1}^{\mathcal{B}}$.
Algorithm 7 Model Predictive Task Sampling
Appendix A Quick Guideline to MPTS
Task episodic learning serves as a cornerstone in developing adaptive models by structuring diverse, context-rich learning experiences. One of the pivotal insights underpinning this process is the neural scaling law, which establishes a relationship between task volume, model complexity, and computational resources, offering a principled insight into training foundation models at a certain budget. Recent viewpoints have also shed light on the importance of task quality 17, 29, 19, 30, 20, 64, 18, 65, prompting innovative data curation strategies to refine datasets for pretraining, meta-training, and post-training. Evidence suggests that carefully curated data can significantly reduce task sampling complexity, decrease computational demands, and enhance robustness against distributional shifts—sometimes achieving these goals simultaneously. Despite these advancements, a practical operation such as Evaluate-Rank-Filter still faces challenges associated with costly evaluations from intensive task queries, computational overhead, and massive annotations. Addressing these bottlenecks remains essential to fully realize the potential of task episodic learning in robust efficient foundation model training.
Computational Complexity Analysis.
The involvement of the risk learner inevitably brings extra computational overhead in optimization. However, the risk learner used in this work is lightweight with the model complexity $\mathcal{O}(|\bm{\phi}|+|\bm{\psi}|)<<\mathcal{O}(|\bm{\theta}|)$. We can roughly estimate these extra computations that arise from the predictive model as $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|)T_{\text{MPTS}}\big{)}$ throughout the training phase. Moreover, the computational and task evaluation complexities of different methods are estimated in Table 2. Compared with DRM, MPTS retains more computational and task efficiency when the filtering ratio $\hat{\alpha}$ is high, and the machine learner $\bm{\theta}$ is largely given similar convergence iteration steps. However, exactly evaluating task difficulties from either losses or gradients brings extra computational and annotation overhead.
Table 2: Computational Complexities using Different Methods. Here, we drop out the ranking or reweighting computational complexity as the model complexity of the machine learner considered in this analysis is major, such as the multimodal foundation models. $T$ refers to the required iteration steps until the convergence for separate methods.
ERM DRM GDRM MPTS (Ours)
computation $\mathcal{O}(|\bm{\theta}|T_{\text{ERM}})$ $\mathcal{O}(\frac{1}{1-\hat{\alpha}}|\bm{\theta}|T_{\text{DRM}})$ $\mathcal{O}(|\bm{\theta}|T_{\text{GDRM}})$ $\mathcal{O}\big{(}(|\bm{\phi}|+|\bm{\psi}|+|\bm{\theta}|)T_{\text{MPTS}}\big{)}$
task eval $\mathcal{O}(\mathcal{B}T_{\text{ERM}})$ $\mathcal{O}(\frac{\mathcal{B}}{1-\hat{\alpha}}T_{\text{DRM}})$ $\mathcal{O}(\mathcal{B}T_{\text{GDRM}})$ $\mathcal{O}(\mathcal{B}T_{\text{MPTS}})$
Choice of Surrogate Models.
Among MPTS’s core components, the risk learner works to predict the adaptation risk values based on historical information and further serves the calculation of acquisition functions. Importantly, this work primarily investigates the feasibility and effectiveness of MPTS and does not impose rigid constraints on the form of the risk learner $p(\ell|\bm{\tau},H_{1:t})$ too much in modeling. The design of this risk learner $p(\ell|\bm{\tau},H_{1:t})$ must meet several criteria: it is tractable in optimization, can process historical risk information, and offers uncertainty in prediction.
A series of candidate probabilistic models exist that probably apply to adaptation risk modeling. One alternative choice can be the Gaussian process 66, which provides an analytical form of the predictive distribution. However, its implementation (i) is less scalable in the case of relatively higher dimensional task identifiers, (ii) holds the cubic runtime complexity in obtaining the predictive covariance matrix, (iii) is sensitive to kernel selection, coupled with limited expressiveness of the Gaussian distribution in learned risk functions. Hence, for simplicity and computational efficiency, we adopt the basic VAE-like model and execute a handful of gradient updates to train the risk learner. We leave more advanced risk learner modeling for future exploration.
Bayesian Optimization for Black-box Functions.
This work relates to active sampling and Bayesian optimization. The purpose of BO 67 is to sequentially find a global optimum of a black-box function $f(\bm{x})$ expensive to evaluate in $\mathcal{S}$, namely $\bm{x}_{*}=\arg\max_{\bm{x}\in\mathcal{S}\subset\mathbb{R}^{d}}f(\bm{x})$.
In each iteration $t=1,\dots,T$, the BO method actively queries $\bm{x}_{t}$ to evaluate $f(\bm{x}_{t})$, yeilding an output $\ell_{t}=f(\bm{x}_{t})+\epsilon$ with a white noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. Due to the high cost of function evaluation, the key to BO is constructing a surrogate model to guide the data point to query. The resulting acquisition function 68 works as an active sampling objective to maximize and obtain the candidate $\bm{x}_{t}$ based on the previous sequence. BO requires limited function evaluations as observations and exploits the correlations in queried data points. These properties make it more theoretically data efficient than random or grid search in seeking the optimal solution 69. This work differs from standard BO as task episodic learning is not the optimal parameter search problem.
Specific Pseudo Algorithms in Considered Scenarios.
The main paper provides the workflow of MPTS in Algorithm 1. For separate scenarios, we attach detailed pseudo algorithms as follows. These illustrated Algorithms are in the context of supervised learning. Regarding RL scenarios, such as meta RL and DR, there is a slight modification for MPTS. As simply picking up worst-case MDPs restricts the task subspace in optimization 33, we adopt the mixture of the identifier subset from the random sampler and the identifier subset from the MPTS sampler. For example, in meta RL, with the pseudo batch size ^$\hat{\mathcal{B}}=1.5\mathcal{B}$, there $1.5\mathcal{B}$ identifier candidates from the random sampler. We retain $0.5\mathcal{B}$ random ones and execute standard MPTS amortized evaluation and acquisition rule to obtain another $0.5\mathcal{B}$ ones from the rest random $\mathcal{B}$ identifiers, formulating the mixed $\mathcal{B}$ task batch for RL training. Such an operation makes RL over the MDP distribution stable in optimization.
Appendix B Research Background
B.1 Any-Shot Cross-Task Generalization
Learning from zero-shot or few-shot examples has been identified as a crucial adaptation capability of the machine learner nowadays 70, 71. This work refers to traditional finetuning as many-shot learning and treats the individual example as each task in the MPTS setup. As finetuning techniques have been widely discussed in the field 61, we skip this part in the background introduction.
Zero-Shot Adaptation.
This assesses the machine learner’s generalization capability when directly deploying in unseen scenarios without the help of a support dataset. Such a cross-task generalization is commonly studied in computer vision 72, and the core of the relevant methods is effective semantic representation either from embedding-based methods 73, 74, 75 or generative-based methods 76, 77. In the era of the foundation models, the pretraining mechanism between multimodality also sometimes empowers the machine learner, such as CLIP 1, with zero-shot capability. When it comes to sequential decision-making, a commonly used technique is called DR 43, 9. In detail, DR places a distribution over environments for the agent to interact, deriving a generalist policy to play quickly in unseen scenarios.
Few-Shot Adaptation.
This examines the machine learner’s capability of resolving unseen tasks from some annotated examples as hints. Meta-learning, as the typical learning paradigm, has gained popularity over the past decade. It achieves few-shot adaptation by leveraging past experience and distilling knowledge to unseen but similar scenarios in a few-shot way 78. In brief, we categorize commonly seen methods into context-based, optimization-based, geometric-based, and others. (i) Formulated in an encoder-decoder structure, the context-based method resembles variational autoencoders and encodes the few-shot information into latent variables or embeddings. Typical ones are neural process families 55, 79, 80, 81, which aim to constitute exchangeable deep stochastic processes with neural networks. (ii) The optimization-based methods, with their versatile nature and ability to enable cross-task skill transfer, have piqued the interest and engagement of researchers in the field. For example, MAML 28, 82, 83, 84 reduces meta-learning to a bi-level optimization in the parameter space, and its extensions have been widely investigated in the field. (iii) The deep metric-based methods 85, 86 attempt to embed tasks into the latent space and are more suitable for few-shot image classification tasks. Besides, there are other families, such as hyper-networks 87, 88, recurrent meta-learning 89, etc.
B.2 Dataset Curation and Task-Level Robustness
The risk minimization principle and the adopted sampling criteria are mainly entangled in optimization. Previously, Section 2.2 summarizes the commonly-used risk minimization principles in task episodic learning. Regardless of the adaptation concept and type, we will overview the existing task sampling strategies according to diverse views.
Data Curation in Typical Methods.
In adaptation learning literature, several task curation techniques are devised to pick up high-quality and difficult yet learnable examples 17, 29, 19, 30, 20, and empirical findings examine their role in yielding robustness enhancement and efficiency gains. There are other coreset methods that find a small portion of the task dataset to represent the full batch example’s utility, e.g., via gradient approximation in optimization 64, 18, 65. The studied framework brings a subproblem to solve, and its primary goal is to achieve data efficiency. Hence, the acquisition strategy used in MPTS can be viewed as the episodic coreset selection for robustness purposes.
Task Distributional Robustness.
The CVaR or expected shortfall 31 is a statistical measure to assess the proportional worst-case performance of some models at certain levels. This is widely adopted in risk-averse applications. As implied in Definition 1, $\text{CVaR}_{\alpha}$ describes the expected risk under the normalized $(1-\alpha)$ proportional tail risk task distribution, and this work specifies the distribution in the task space. Meanwhile, the normalized tail task distribution $p_{\alpha}(\tau;\bm{\theta})$ can be viewed as a shifted result from the initial task distribution $p(\tau)$; hence, such a measure provides robustness quantification in the presence of task distribution shifts 16, 46, 33.
Definition 1 (Conditional Value-at-Risk (CVaR) 31)
Given the machine learner parameter $\bm{\theta}$, we denote the task specific random variable by $\ell_{i}:=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta})$. Throughout the task space $\mathcal{T}$, let the cumulative risk distribution and the quantile of risk values respectively be $F(\ell)$ and $\ell^{\alpha}=\min_{\ell}\{\ell|F(\ell)\geq\alpha\}$. Then the CVaR at $\alpha$-robustness level can be estimated as:
[…]
where we define the normalized cumulative distribution of task risk values by:
$\begin{split}F^{\alpha}(\ell;\bm{\theta})=\begin{cases}0,&l<\ell^{\alpha}\\ \frac{F(\ell;\bm{\theta})-\alpha}{1-\alpha},&l\geq\ell^{\alpha}.\end{cases}\end{split}$ (21)
And this induce the tail risk task distribution denoted by $p_{\alpha}(\tau;\bm{\theta})$.
Another strategy to evaluate the machine learner’s robustness is the performance in OOD tasks. This refers to the case when the training and the testing task distributions are different. Particularly, in DR and prompt-tuning scenarios, we also use the OOD tasks that never appear in the training task distribution to test the trained policy, and this setup corresponds to domain generalization, a type of substantial distribution shift 11.
B.3 Baseline Details
GDRM 35.
As implied in the main paper, this can be interpreted as a min-max optimization problem. The basic idea of increasing the machine learner’s robustness lies in assigning more probability mass to worst cases in a reweighted manner. That means in each iteration with the best selected $p_{\hat{g}}(\tau)$, the optimization problem is reduced as
$\begin{split}\min_{\bm{\theta}\in\bm{\Theta}}\mathbb{E}_{p_{\hat{g}}(\tau)}\Big{[}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\Big{]}=\mathbb{E}_{p(\tau)}\left[\frac{p_{\hat{g}}(\tau)}{p(\tau)}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\right],\end{split}$ (22)
where we use $\omega(\tau)=\frac{p_{\hat{g}}(\tau)}{p(\tau)}$ to denote the weight.
Given a fixed number of tasks, GDRM will heuristically or dynamically group them into clusters and then perform a reweighting mechanism according to the evaluated risk. In task episodic learning, there is no task grouping operation as the task batch is reset after each iteration 16, 46. And the default computation of task-specific weigths is $\omega(\tau_{i})=\frac{\exp(\eta\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}))}{\sum_{b=1}^{\mathcal{B}}\exp(\eta\ell(\mathcal{D}_{\tau_{b}}^{Q},\mathcal{D}_{\tau_{b}}^{S};\bm{\theta}))}$, where $\eta$ is the temperature parameter and $\{\bm{\tau}_{b}\}_{b=1}^{\mathcal{B}}$ is the identifier of the task batch. The implementation detail can be found in https://github.com/kohpangwei/group_DRO.
DRM 16, 46.
Similar to GDRM, the optimization objective $\mathbb{E}_{p_{\alpha}(\tau;\bm{\theta})}\Big{[}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})\Big{]}$ also defines the task distribution with constraints. As $p_{\alpha}(\tau;\bm{\theta})$ is the machine learner’ parameter $\bm{\theta}$dependent, a commonly used practical implementation is to evaluate the performance of the machine learner, rank the task-specific adaptation risk value, and then select $(1-\alpha)$ proportional worst cases to optimize.
In this work, we retain the setup in 16 and pick up the $-\mathcal{B}$ in optimization. This implies that the actual task batch to evaluate is $\frac{\mathcal{B}}{1-\alpha}$. And for fair comparison with MPTS and light computations, we retain the Monte Carlo estimator for the risk quantile in implementation, which is same as 16. To ensure stable training, for all benchmarks, we keep the actual task batch ^$\hat{\mathcal{B}}=2\mathcal{B}$ to evaluate and discard the easiest half before the machine learner’s optimization.
As revealed in works 90, 38, 16, the heuristic operation as the Evaluate-Rank-Filter or reweighting mechanism in GDRM is widely adopted for approximate optimization. For example, in task robust meta-learning scenarios, the prerequisite step in DR-MAML 16 is to execute gradient updates in the inner loop for all candidate tasks and then screen the tail task subset to meta-optimize according to the evaluation results.
Appendix C Task Formulations and Identifiers
Here we refer to the variables that sufficiently configure a task as the task identifier $\bm{\tau}$. In other literature work, these task identifiers can be viewed as the task representations in a lower dimensional space. To clarify these concepts, we provide more explanations in specific scenarios.
C.1 Tasks with Explicit Identifiers
K-shot Sinusoid Regression.
In $K$-shot sinusoid regression 28, meta learners aim at quickly adapting the model to an unseen function $f(x)=a\sin(x-b)$ with the help of $K$ data points randomly sampled from the function. This case treats the amplitude and phase variables $(a,b)$ as the task identifier to configure the task. And the task distribution is induced by the uniform distribution over the task identifier.
Meta Reinforcement Learning.
Here, we take the ReacherPos task as an example. The goal of the robot arm is to reach an unobserved target location $[x_{1},x_{2}]$. The end-effector position of the robot arm is initialized randomly, and the step-wise reward corresponds to the feedback to the agent after each move based on its distance to the target location. As the task distribution is specified by a uniform distribution over the target location, $\bm{\tau}=[x_{1},x_{2}]$ can be viewed as the task identifier. Similarly, we vary physics parameters in simulators to generate diverse MDPs. This constitutes different meta RL benchmarks.
Domain Randomization.
DR is a promising paradigm to achieve zero-shot adaptation in unseen scenarios, which is widely adopted in robotics 9 and computer vision 43. The basic idea is to train the machine learner in a wide variety of environments or deployment scenarios and then directly apply the learned model to new ones.
Table 3: Benchmarks with Explicit Task Identifiers. Here, we list the detail information about the task identifier to induce the task distribution.
     Benchmarks      Identifier Meaning      Identifier Range
     K-shot sinusoid regression      amplitude and phase $(a,b)$      $[0.1,5.0]\times[0,\pi]$
     Meta-RL: HalfCheetahMassVel      mass and velocity $(m,v)$      $[0.75,1.25]\times[0,2.0]$
     Meta-RL: HalfCheetahVel      velocity $v$      $[0,2.0]$
     Meta-RL: ReacherPos      goal location $(x_{1},x_{2})$      $[-0.2,0.2]\times[-0.2,0.2]$
     Meta-RL: Walker2dMassVel      mass and velocity $(m,v)$      $[0.75,1.25]\times[0,2.0]$
     Meta-RL: Walker2dVel      velocity $v$      $[0,2.0]$
     DR: LunarLander      main engine strength $s$      $[4,20]$
     DR: ErgoReacher      joint damping $d$ and max torque $t$ ($\times$4 joints)      $[0.1,2.0]\times[2,20]$
As noted in the main paper, we suppose that the task identifier contains semantics that reflects the difficulty of tasks to resolve and the adaptation risk function is smooth with respect to the identifier. In total, we summarize these bechmarks with explicit task identifiers in Table 3.
C.2 Tasks with Implicit Identifiers
As previously mentioned, we assume the existence of a statistical correlation between task identifiers and adaptation risk values given a specific adaptive machine learner. This implies that the task identifier preserves precise semantics about the task information. These provide the basis for establishing the risk learner from the coupled dataset $\{[\bm{\tau}_{i},\ell_{i}]\}_{i=1}^{\mathcal{B}}$.
Nevertheless, in several scenarios, it is intractable to access the explicit task identifier. For example, in few-shot image classification, the task information is just the coupled support and query dataset $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{S}\cup\mathcal{D}_{\tau}^{Q}$. Similarly, in PEFT for LLMs, the task can be in the form of the QA pair $\mathcal{D}_{\tau}=\mathcal{D}_{\tau}^{Q}$. There is no explicit representation method, such as $\bm{\tau}$, for these tasks, which brings difficulty in building up the risk learner. Retaining the prior notation, the episodic task batch can be written as $\hat{H}_{t}=\left\{\bm{\theta}_{t},\left(\bm{\tau}_{t,i},\mathcal{D}_{\tau_{t,i}},\ell_{t,i}\right)\right\}_{i=1}^{\mathcal{B}}$, where $\bm{\tau}$ of our interest is unobservable. Some experiments in this work, such as few-shot image classification and PEFT, encounter such embarrassment.
Task Representation through Identifier Inference.
To scale our approach under these circumstances, we propose an alternative candidate schema as the complementary. The probabilistic relationship between variables is depicted in Fig. 2. We consider obtaining the implicit identifier through inference from the task dataset. To do so, we include additional module $f_{\xi}$ with $\xi\in\Xi$ to embed $\mathcal{D}_{\tau}^{S}$ and $\mathcal{D}_{\tau}^{Q}$ and further induce a vector $\bm{\tau}=f_{\xi}(\mathcal{D}_{\tau}^{S},\mathcal{D}_{\tau}^{Q})$ as the approximate task identifier. These operations imply seeking the appropriate inference module directly influences the risk learner’s performance.
Fortunately, there exist pretrained models that enable the task representation to be generalizable to downstream tasks. For example, in the N-way K-shot image classification, the task is in the form of support image-label pairs and the query images and the goal is to assign labels to the query images from the support dataset. With the help of CLIP models 1, for a fixed task in the form of $\mathcal{D}_{\tau}$, we can access a N vectors $\{\bm{z}_{i}\}_{i=1}^{N}$ by inputing the set of text-form classes $\{\mathcal{C}_{i}\}_{i=1}^{N}$ extracted from the support dataset $\mathcal{D}_{\tau}^{S}$, i.e., $\text{CLIP}(\{\mathcal{C}_{i}\}_{i=1}^{N})=[\text{CLIP}_{\text{text}}(\mathcal{C}_{1}),\dots,\text{CLIP}_{\text{text}}(\mathcal{C}_{K})]:=\bm{\tau}$. As a result, we can obtain $H_{t}=\{[\bm{\tau}_{t,i},\ell_{t,i}]\}_{i=1}^{\mathcal{B}}$ conditioned on $\bm{\theta}_{t}$ for feasible task risk functional prosterior inference. This helps our approach to circumvent the unavailability of exact task identifiers. And it is plausible for the risk predictive model to optimize in learning $p(\ell|\bm{\tau},H_{1:t})$. It is worth noting that this case still prefers lightweight models for identifier inference, and the text encoder of CLIP well satisfies this requirement and can be used in the N-way K-shot image classification. Details on specific task identifier inference modules can be found in Section E and F.
Appendix D Auto-Encoding Adaptation Risk through Streaming VI
Note that the basis of MPTS is to establish the bridge between the task identifier and the adaptation risk value over the course of the machine learner’s optimization. This implies we are seeking a lightweight stochastic risk function in Definition 2 to approximate the posterior $p(\ell|\bm{\tau},H_{1:t})$ in the task space.
Definition 2 (Stochastic Risk Function)
Let $\mathfrak{X}$ denote the index set’s Cartesian product with the task identifier’s dimension $\bm{\tau}\in\mathbb{N}^{d}$. For any $k\in\mathbb{N}$ and finite index sequence $\bm{\tau}_{1},\dots,\bm{\tau}_{k}\in\mathfrak{X}$, we write some probability measure over $\mathbb{R}^{k}$ as $\nu_{(\bm{\tau}_{1},\dots,\bm{\tau}_{k})}$. By introducing the probability space $(\Omega_{\tau},\mathcal{F}_{\bm{\theta}},\mathcal{P})$ and $\forall\theta\in\Theta$, we can induce a stochastic function $\mathcal{F}_{\bm{\theta}}:\mathcal{T}\times\Omega_{\tau}\mapsto\mathbb{R}^{k}$, so that $\nu_{(\bm{\tau}_{1},\dots,\bm{\tau}_{k})}(C_{1}\times\dots\times C_{k})=\mathcal{P}(\mathcal{F}_{\bm{\theta}}(\bm{\tau}_{1})\in C_{1},\dots,\mathcal{F}_{\bm{\theta}}(\bm{\tau}_{k})\in C_{k})$ $\forall\bm{\tau}_{i}\in\mathfrak{X}$ and $C_{i}\in\mathbb{R}$.
This section details steps in auto-encoding historical task risk information, parameterizing variational distributions, deriving the approximate optimization objective, and estimating the stochastic gradients of parameters.
D.1 Neural Modules to Parameterize Distributions
Here, we detail the neural modules to parameterize the distributions of interest. For the approximate posterior $q_{\bm{\phi}}(\bm{z}_{t}|H_{t})$ and conditional prior $p(\bm{z}_{t}|H_{1:t-1})$, the inputs of the module are a set of task risk pairs. The neural module requires the permutation invariance w.r.t. the order of the data points in the set $H_{t}$ or $H_{1:t-1}$ in Definition 3. Hence, we adopt the DeepSet style neural network 56 to process the collected $H_{t}$ or $H_{1:t-1}$.
Definition 3 (Permutation Invariant Function)
With an $n$-element permutation group $\mathcal{S}_{n}$, the operator $g\in\mathcal{S}_{n}$ maps the order set to itself:
$g:[1,2,\dots,n]\mapsto[g_{1},g_{2},\dots,g_{n}].$ (23)
Then the function $\Phi$ is called permutation invariant if for any set of data points $\bm{x}_{1},\dots,\bm{x}_{n}$, the following condition holds:
$\Phi(g\circ[\bm{x}_{1},\dots,\bm{x}_{n}])=\Phi([\bm{x}_{g_{1}},\dots,\bm{x}_{g_{n}}])=\Phi([\bm{x}_{1},\dots,\bm{x}_{n}])\quad\forall g\in\mathcal{S}_{n}.$ (24)
For example, we denote the neural network parameters by $\bm{\phi}=\{\bm{\phi}_{1},\bm{\phi}_{2,1},\bm{\phi}_{2,2}\}$ together with a mean pooling operator $\oplus$, we can have:
$\begin{split}\bm{r}_{i}=h_{\bm{\phi}_{1}}(\bm{\tau}_{k,i},\ell_{k,i})\ \forall i\in\{1,\dots,\mathcal{B}\},\quad\bar{\bm{r}}=\oplus_{i=1}^{\mathcal{B}}\bm{r}_{i},\quad\bm{\mu}_{\bm{\phi}}=h_{\bm{\phi}_{2,1}}(\bar{\bm{r}})\ \text{and}\ \bm{\Sigma}_{\bm{\phi}}=h_{\bm{\phi}_{2,2}}(\bar{\bm{r}}),\end{split}$ (25)
where the output corresponds to $q_{\bm{\phi}}(\bm{z}_{t}|H_{t})=\mathcal{N}(\bm{\mu}_{\bm{\phi}},\bm{\Sigma}_{\bm{\phi}})$ (see Fig. 8 for details).
Connections with the NP Family. Regarding the task risk functional posterior inference module, this work has a close connection with the NP family 81, 55, 91, 92, 93, 94, 79. Both handle the set data points in probabilistic inference.
Also, note that this work is significantly distinguished from the NP in terms of research motivations and learning setups. The NP mostly copes with the stationary stochastic process and approximates the underlying functional prior with the help of a context-based latent variable. Our setup considers learning in the non-stationary case, where histories arrive sequentially and depend on the updated machine learner. Hence, we adopt the streaming VI 40 and capture the task risk value in a posterior inference manner.
D.2 Formulation of ELBO & Stochastic Gradient Estimates
Figure 8: The Encoder-Decoder Neural Network to Paramterize the Risk Learner.
Unlike previous risk minimization principles in task episodic learning, ours include an additional risk predictive module, which guides the task batch sampling. Importantly, we use the latent variable to summarize the historical information information and quantify uncertainty in predicting task-specific adaptation risk. The following details the steps.
[…]
$\geq\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\Big{[}\ln p_{\bm{\psi}}(H_{t}|\bm{z}_{t})\Big{]}-D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel p(\bm{z}_{t}|H_{1:t-1})\Big{]}:=\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})$ (26c)
Then, we can rewrite the ELBO with the help of reparameterization trick 25 in Eq. (27).
$\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\Big{[}\ln p_{\bm{\psi}}(H_{t}|\bm{z}_{t})\Big{]}-D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel p(\bm{z}_{t}|H_{1:t-1})\Big{]}$ (27a)
$=\mathbb{E}_{p(\bm{\epsilon})}\Big{[}\ln p_{\bm{\psi}}(H_{t}|g_{\bm{\phi}}(\bm{\epsilon},H_{t}))\Big{]}-D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel p(\bm{z}_{t}|H_{1:t-1})\Big{]}$ (27b)
$\approx\ln p_{\bm{\psi}}(H_{t}|g_{\bm{\phi}}(\bm{\epsilon},H_{t}))-D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel p(\bm{z}_{t}|H_{1:t-1})\Big{]},\quad\text{with}\quad\bm{\epsilon}\sim\mathcal{N}(0,\bm{I}_{d})$ (27c)
Moreover, we estimate the stochastic gradients w.r.t. all model parameters based on the reparameterized latent variable distribution.
$\nabla_{\bm{\phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})\approx\nabla_{\bm{\phi}}\ln p_{\bm{\psi}}(H_{t}|g_{\bm{\phi}}(\bm{\epsilon},H_{t}))-\frac{1}{2}\nabla_{\bm{\phi}}\Big{(}\text{Tr}(\hat{\bm{\Sigma}}^{-1}\bm{\Sigma}_{\bm{\phi}})+(\hat{\bm{\mu}}-\bm{\mu}_{\bm{\phi}})^{T}\hat{\bm{\Sigma}}(\hat{\bm{\mu}}-\bm{\mu}_{\bm{\phi}})-\ln(\det\bm{\Sigma}_{\bm{\phi}})\Big{)}$ (28a)
$\text{with}\quad q_{\bm{\phi}}(\bm{z}_{t}|H_{t})=\mathcal{N}(\bm{\mu}_{\bm{\phi}},\bm{\Sigma}_{\bm{\phi}})\ \text{and}\ p(\bm{z}_{t}|H_{1:t-1})=\mathcal{N}(\hat{\bm{\mu}},\hat{\bm{\Sigma}})$ (28b)
$\nabla_{\bm{\psi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi})\approx\nabla_{\bm{\psi}}\ln p_{\bm{\psi}}(H_{t}|g_{\bm{\phi}}(\bm{\epsilon},H_{t}))$ (28c)
As illustrated in Eq. (28), one stochastic forward pass is required for gradient estimates in the training process. For flexible implementation, we adopt a $\beta$-VAE strategy to turn Eq. (27) into
$\begin{split}\max_{\bm{\psi}\in\bm{\Psi},\bm{\phi}\in\bm{\Phi}}\mathcal{G}_{\text{ELBO}}(\bm{\psi},\bm{\phi}):=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}_{t}|H_{t})}\left[\sum_{i=1}^{\mathcal{B}}\ln p_{\bm{\psi}}(\ell_{t,i}|\bm{\tau}_{t,i},\bm{z}_{t})\right]-\beta D_{KL}\Big{[}q_{\bm{\phi}}(\bm{z}_{t}|H_{t})\parallel q_{\bar{\bm{\phi}}}(\bm{z}_{t}|H_{t-1})\Big{]}\end{split}$ (29)
D.3 Theoretical Guarantee for Task Difficulties’ Scoring with Posterior Inference
Assumption 1 (Lipschitz Continuity) We assume the adaptation risk function $\ell(\cdot;\bm{\theta})$ reserves the Lipschitz continuity w.r.t. $\bm{\theta}$ and $\bm{\tau}$, i.e.,
$\begin{split}|\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})-\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}^{\prime})|\leq\beta_{1}||\bm{\theta}-\bm{\theta}^{\prime}||\quad\text{and}\quad|\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})-\ell(\mathcal{D}_{\tau^{\prime}}^{Q},\mathcal{D}_{\tau^{\prime}}^{S};\bm{\theta})|\leq\beta_{2}||\bm{\tau}-\bm{\tau}^{\prime}||,\end{split}$ (30)
where $\forall\{\bm{\theta},\bm{\theta}^{\prime}\}\in\bm{\Theta}$ and $\forall\{\bm{\tau},\bm{\tau}^{\prime}\}\in\mathcal{T}$ with Lipschitz constants $\beta_{1}$ and $\beta_{2}$.
Assumption 2 (Bounded Sample Gradient) We assume the norm of the adaptation risk function’s gradient $\nabla\ell(\cdot;\bm{\theta}_{t})$ is bounded:
$\sup_{\tau\in\mathcal{T}}||\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta}_{t})||_{2}<G_{t},$ (31)
where $G_{t}$ is a positive constant.
Assumption 3 (Sub-Gaussian Stochastic Gradient) The stochastic gradient $\tilde{\bm{g}}:=\bm{g}+\bm{\epsilon}$ for the machine learner’s adaptation at $t$-th iteration is $\sigma$-sub-Gaussian, which means:
$\begin{split}\mathbb{E}\left[\exp{(\eta\bm{v}^{T}\bm{\epsilon})}\right]\leq\exp{\left(\frac{\eta^{2}\sigma^{2}||\bm{v}||_{2}^{2}}{2}\right)}\quad\forall\eta\in\mathbb{R}\ \text{and}\ \bm{v}\in\mathbb{R}^{d},\end{split}$ (32)
where $\mathbb{E}[\tilde{\bm{g}}]=\bm{g}$, $\mathbb{E}[||\tilde{\bm{g}}-\bm{g}||_{2}^{2}]\leq\sigma^{2}$ and $\sigma\in\mathbb{R}^{+}$.
Given the Assumption 3 and the Chernoff bound 95, we can have the concentration inequality as:
$\begin{split}\mathbb{P}(||\tilde{\bm{g}}-\bm{g}||_{2}\geq t)\leq 2\exp{(-\frac{t^{2}}{2\sigma^{2}})}\quad\forall t\in\mathbb{R}.\end{split}$ (33)
Theorem 1 (Provably Approximately Invariant Task Difficulties) Given arbitrary $K$ data points $\{(\bm{\tau}_{i},\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})\}_{i=1}^{K}$, the adaptation gradient $\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})$ as a $\sigma$-sub-Gaussian random variable and $\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})$, we denote the relative difficulty via the difference $\Delta_{ij}(\bm{\theta}_{t+1})=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t+1})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t+1})$ and $\Delta_{ij}(\bm{\theta}_{t})=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})$ between $t$-th and $(t+1)$-th iterations, and the gradient difference as $\bm{v}_{ij}:=\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})$.
Under Assumption 1/2/3, the set of rank-preserving variable $E_{ij}:=\mathds{1}\left[\text{sign}(\Delta_{ij}(\bm{\theta}_{t+1}))=\text{sign}(\Delta_{ij}(\bm{\theta}_{t}))\right]$ satisfies the probability inequality:
$\mathbb{P}(\bigcap_{i<j}E_{ij})\geq 1-\xi,$
when $\eta\leq\frac{\delta_{t}}{2G_{t}M_{t}+\sqrt{8\sigma^{2}G_{t}^{2}\ln\left(\frac{K(K-1)}{2\xi}\right)}}$ with $G_{t}$ in Assumption 2, $\delta_{t}:=\min_{i\neq j}|\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})|\in\mathbb{R}^{+}$, the stochastic gradient norm $M_{t}:=||\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})||_{2}$.
The purpose of this part is to uncover the mechanism of the risk learner in amortized evaluation of adaptation risk values and scoring the difficulty of tasks. The function of the risk learner relies on Assumptions 1/2/3 and the posterior inference $p(\ell|\tau,H_{1:t};\bm{\theta}_{t})$ from the historical risk information $H_{1:t}$. The foundation of predicting the outcome of optimization in a rough granularity lies in the Theorem 1, and we detail the proof of such a theorem as below.
①. Any-Shot Adaptation After One-step Gradient Descent.
Here, we consider a set of data points for the risk learner $\{(\bm{\tau}_{i},\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})\}_{i=1}^{K}$ under an arbitrary fixed machine learner $\bm{\theta}_{t}$, where tasks in the set $\{\tau_{i}\}_{i=1}^{K}$ are randomly sampled from $p(\tau)$. Without loss of generality, we can assume that the adaptation risk values satisfy a rank ordering:
$\ell(\mathcal{D}_{\tau_{1}}^{Q},\mathcal{D}_{\tau_{1}}^{S};\bm{\theta}_{t})\geq\ell(\mathcal{D}_{\tau_{2}}^{Q},\mathcal{D}_{\tau_{2}}^{S};\bm{\theta}_{t})\geq\cdots\geq\ell(\mathcal{D}_{\tau_{K}}^{Q},\mathcal{D}_{\tau_{K}}^{S};\bm{\theta}_{t}).$ (34)
The gradient descent as fast adaptation:
$\begin{split}\bm{\theta}_{t+1}=\bm{\theta}_{t}-\eta\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t}).\end{split}$ (35)
After the above operator, we can obtain another set of data points for the updated risk learner $\{(\bm{\tau}_{i},\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t+1})\}_{i=1}^{K}$.
②. Changes of Adaptation Risk Values and Pairwise Ranks.
Based on the Assumption 1, we can perform local approximation over $\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta})$ with the help of first-order Talor expansion w.r.t. the $\bm{\theta}_{t}$:
$\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t+1})=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\eta\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})^{T}\mathcal{\bm{L}}(\bm{\theta}_{t})+\mathcal{O}(\eta^{2}||\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})||_{2}^{2})$ (36a)
$\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t+1})\approx\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})^{T}\mathcal{\bm{L}}(\bm{\theta}_{t})\quad\forall i\in\{1,2,\dots,K\}.$ (36b)
One straightforward way to assess the task difficulty is to compare arbitrary paired tasks $\{\bm{\tau}_{i},\bm{\tau}_{j}\}$’s adaptation risk values $\{\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t}),\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})\}$ with $i<j$. Then, we can estimate the relative difficulty via the difference as:
$\begin{split}\Delta_{ij}(\bm{\theta}_{t+1})\approx\Delta_{ij}(\bm{\theta}_{t})-\eta\left(\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})\right)^{T}\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t}),\end{split}$ (37)
where we denote the relative difficulty via the difference as $\Delta_{ij}(\bm{\theta}_{t+1})=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t+1})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t+1})$ and $\Delta_{ij}(\bm{\theta}_{t})=\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})$ between $t$-th and $(t+1)$-th iterations. As $\Delta_{ij}(\bm{\theta}_{t})$ is positive, one feasible condition for $\Delta_{ij}(\bm{\theta}_{t+1})\in\mathbb{R}^{+}$ is:
$\Delta_{ij}(\bm{\theta}_{t+1})\approx\Delta_{ij}(\bm{\theta}_{t})-\eta\left(\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})\right)^{T}\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})$ (38a)
$\geq\Delta_{ij}(\bm{\theta}_{t})-2\eta G_{t}M>0,\quad\Rightarrow\eta<\frac{\Delta_{ij}(\bm{\theta}_{t})}{2G_{t}M_{t}}\quad\text{with}\quad M_{t}:=||\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})||_{2}.$ (38b)
The above implies that when the learning rate $\eta$ in gradient step is smaller enough, the relative difficulty between the task $i$ and $j$ can be preserved after the machine learner’s update with the Assumption 2.
③. Probabilistic Inequality with a Nearly Invariant Ranking Guarantee.
In practice, the stochastic gradient descent is performed, which means the gradient is a random variable $\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})=\bm{g}_{t}+\bm{\epsilon}$ with $\mathbb{E}[\bm{\epsilon}]=0,\mathbb{E}[||\bm{\epsilon}||_{2}^{2}]<\sigma^{2}$ and $\bm{g}_{t}=\mathbb{E}\left[\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})\right]$. Meanwhile, we denote the gradient difference by $\bm{v}_{ij}:=\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})$, which leads to:
$\begin{split}||\bm{v}_{ij}||_{2}\leq 2\sup_{\tau\in\mathcal{T}}||\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})||_{2}<2G_{t},\end{split}$ (39)
according to the Assumption 2. Another variable is introduced as the minimum separation between arbitrary paired adaptation risk values:
$\begin{split}\delta_{t}:=\min_{i\neq j}|\ell(\mathcal{D}_{\tau_{i}}^{Q},\mathcal{D}_{\tau_{i}}^{S};\bm{\theta}_{t})-\ell(\mathcal{D}_{\tau_{j}}^{Q},\mathcal{D}_{\tau_{j}}^{S};\bm{\theta}_{t})|\in\mathbb{R}^{+}.\end{split}$ (40)
Still, to make sure the invariant rank, one necessary condition can be:
$\begin{split}\Delta_{ij}(\bm{\theta}_{t})-\eta\bm{v}_{ij}^{T}\bm{g}_{t}\geq 0\end{split}$ (41)
And the above inequality reasonably holds when $\eta\bm{v}_{ij}^{T}\bm{g}_{t}<\delta_{t}$. Here, we define the random event $E_{ij}:=\mathds{1}\left[\text{sign}(\Delta_{ij}(\bm{\theta}_{t+1}))=\text{sign}(\Delta_{ij}(\bm{\theta}_{t}))\right]$ from the task pair together with $E_{ij}^{c}:=\mathds{1}\left[\text{sign}(\Delta_{ij}(\bm{\theta}_{t+1}))\neq\text{sign}(\Delta_{ij}(\bm{\theta}_{t}))\right]$. With the help of $\sigma$-sub-Gaussain property in Assumption 3, we can bound the case of the rank flipping as (note some critical conditions that $\bm{v}_{ij}^{T}\bm{g}_{t}\in\mathbb{R}^{+}$ and $\eta\bm{v}_{ij}^{T}\bm{g}_{t}<\delta_{t}$ as the learning rate $\eta$ can be typically smaller enough):
$\mathbb{P}(E_{ij}^{c})=\mathbb{P}(\eta\bm{v}_{ij}^{T}\nabla_{\bm{\theta}}\mathcal{\bm{L}}(\bm{\theta}_{t})\geq\delta_{t})\leq\exp{\left(-\frac{(\delta_{t}-\eta\bm{v}_{ij}^{T}\bm{g}_{t})^{2}}{2\eta^{2}\sigma^{2}||\bm{v}_{ij}||_{2}^{2}}\right)}<\exp{\left(-\frac{(\delta_{t}-2\eta G_{t}M_{t})^{2}}{8\eta^{2}\sigma^{2}G_{t}^{2}}\right)}$ (42a)
$\mathbb{P}(\bigcup_{i<j}E_{ij}^{c})\leq\sum_{i<j}\mathbb{P}(E_{ij}^{c})\leq\frac{K(K-1)}{2}\exp{\left(-\frac{(\delta_{t}-2\eta G_{t}M_{t})^{2}}{8\eta^{2}\sigma^{2}G_{t}^{2}}\right)}$ (42b)
$\mathbb{P}(\bigcap_{i<j}E_{ij})=1-\mathbb{P}(\bigcup_{i<j}E_{ij}^{c})\geq 1-\frac{K(K-1)}{2}\exp{\left(-\frac{(\delta_{t}-2\eta G_{t}M_{t})^{2}}{8\eta^{2}\sigma^{2}G_{t}^{2}}\right)}\geq 1-\xi.$ (42c)
The condition for the above inequality holds is $\eta\leq\frac{\delta_{t}}{2G_{t}M_{t}+\sqrt{8\sigma^{2}G_{t}^{2}\ln\left(\frac{K(K-1)}{2\xi}\right)}}$. With the above steps ①-③ and corresponding conditions, we complete the proof.
Appendix E Prompt-based Few-shot Image Classification
We adopt the standard few-shot image classification setting 28, 85, where tasks are constructed using the N-way K-shot paradigm for both meta-training and meta-testing. Each task comprises support and query sets. The support set contains K examples for each of the N classes, while the query set includes 15 examples per class. During meta-training, labels for both support and query data are accessible to the adaptive machine learner. During meta-testing, the query dataset’s labels are to be predicted given the labeled support dataset. The class categories of task datasets in the meta-training and meta-testing do not overlap. In experiments, we specifically consider a 5-way 1-shot image classificationconfiguration. During meta-training, we set the task batch for ERM and DRM as $\mathcal{B}=4$ (For implementation simplicity, the data loader samples $8$ tasks and then randomly keeps half without ranking to optimize). The task batch for DRM is $\mathcal{B}=8$ before the filtering operation; DRM filters half to optimization. Similarly, that for MPTS is $\mathcal{B}=8$ and only $4$ tasks are screened to optimize.
To enable few-shot learning by prompt-tuning, we integrate the multimodal prompt learning methods MaPLe 44 and prototypical network (ProtoNet) 85. MaPLe operates on the CLIP model 1, capturing multimodal prompts to refine both visual and textual feature representations with frozen CLIP’s parameters. These refined textual features serve as classifiers for predicting refined image features. In parallel, ProtoNet is utilized to derive class-specific visual embeddings from the support set, which assist in distinguishing query samples.
To utilize both MaPLe and ProtoNet, we construct classifiers based on both textual features and visual embeddings. Predictions for query samples are generated by combining the classifiers through a weighted sum. This combination strategy is employed during meta-training to optimize the multimodal prompts. Meanwhile, the CLIP model’s parameters remain frozen throughout optimization. During meta-testing, these trained prompts are adopted to create textual and visual classifiers and process query image features. Final predictions for each query sample are made using the same weighted combination approach as in meta-training.
In mathematics, we can characterize the mentioned pipeline as:
[…]
Classification Likelihood from the Query Dataset: (45)
$\quad p_{\bm{\theta}}(y=k|\bm{x},\bm{u})=\lambda_{1}\frac{\exp(-d(f_{\bm{\theta}_{i}}(\bm{x},\bm{u}),\bm{c}_{k}))}{\sum_{k^{\prime}}\exp(-d(f_{\bm{\theta}_{i}}(\bm{x},\bm{u}),\bm{c}_{k^{\prime}}))}$ $+\lambda_{2}\frac{\exp(-d(f_{\bm{\theta}_{i}}(\bm{x},\bm{u}),\bm{t}_{k}))}{\sum_{k^{\prime}}\exp(-d(f_{\bm{\theta}_{i}}(\bm{x},\bm{u}),\bm{t}_{k^{\prime}}))},$
where $\bm{\theta}_{t}$ and $\bm{\theta}_{i}$ denote the textual and visual encoders of the CLIP model, respectively. $\bm{l}_{k}$ and $\bm{u}$ respectively denote the textual descriptions of category k and the multimodal prompts. $(\bm{x}_{i},y_{i})\in S_{k}$ denotes the images and labels of category k in the support dataset $\mathcal{D}_{\tau}^{S}$. Once the textual classifier $\bm{t}$ and support visual classifier $\bm{c}$ are obtained, we predict the query sample $\bm{x}$ by the classifiers with hyperparameters $\lambda_{1}=0.25$ and $\lambda_{2}=1.0$.
Appendix F Backbone Methods & Experimental Details in Any-Shot Learning
F.1 MAML
In sinusoid regression and Meta-RL, MAML is used as the backbone algorithm. As previously discussed, MAML is widely applied in solving few-shot learning tasks. In mathematics, its optimization objective can be characterized as:
$\begin{split}\min_{\bm{\theta}\in\bm{\Theta}}\mathbb{E}_{p(\tau)}\left[\ell(\mathcal{D}_{\tau}^{Q};\bm{\theta}-\lambda\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{S};\bm{\theta}))\right],\end{split}$ (46)
where the term inside the bracket specifies the adaptation risk $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})$, and $\bm{\theta}-\lambda\nabla_{\bm{\theta}}\ell(\mathcal{D}_{\tau}^{S};\bm{\theta})$ denotes the gradient update with the learning rate $\lambda$ as fast adaptation to the task $\bm{\tau}$. After meta-training, we can access the meta initialization $\bm{\theta}$ that generalizes across the task space.
When it comes to reinforcement learning scenarios, $\mathcal{D}_{\tau}$ corresponds to episodic returns collected from MDPs with either the meta policy or the fast adapted policy. To ensure enough coverage of task space, we adopt a mixture strategy of MPTS and random sampling as an empirical regularizer in all RL scenarios, which is similar to work 33.
F.2 DR
Robotic DR refers to the setup that trains the agent in a collection of environments to obtain a generalizable policy. The diversity of environments tends to increase the robustness of policies in deployment. Such a setup does not require few-shot episodes in unseen but similar environments. In mathematics, we can express the optimization objective as:
$\begin{split}\max_{\bm{\theta}\in\bm{\Theta}}\mathcal{J}(\bm{\theta}):=\mathbb{E}_{\pi_{\bm{\theta}}}\mathbb{E}_{p(\tau)}\left[\sum_{t=0}^{H}\gamma^{t}r_{t}\right]\end{split}$ (47)
where $p(\tau)$ defines the distribution over MDPs, and $\{r_{t}\}_{t=0}^{H}$ is the episodic stepwise reward after interacting with a specific MDP with $H$ as the horizon. Once finishing the optimization of Eq. (47), we can access the policy $\pi_{\bm{\theta}}$ as the zero-shot decision-maker in new environments. In this case, the adaptation risk can be in the form $\ell(\mathcal{D}_{\tau}^{Q},\mathcal{D}_{\tau}^{S};\bm{\theta})=-\sum_{t=0}^{H}\gamma^{t}r_{t}$.
Remember that MDP distribution $p(\tau)$ is mostly induced by physical parameters, e.g., mass, gravity, friction, etc., or the reward functions. In each training iteration, the machine learner resamples a batch of MDPs and gets the shared policy to interact with them to collect episodes. Consequently, the query dataset contains the episodes collected with no support dataset. Overall, the policy optimization follows the standard TD3 algorithm 45 due to its sample efficiency and stability.
F.3 Multi-Modal Prompt Learning
Multi-modal prompt learning is based on the backbone of the prompt tuning method MaPLe 44, which we use on both few-shot and many-shot classifications.
Few-shot classification.
The few-shot prompt learning refers to the common few-shot classification setting 28, 85. We integrate the MaPLe backbone with the ProtoNet 85 to fully utilize the support sets in few-shot learning. As illustrated in Section E, we generate the model prediction using both the textual classifiers from the CLIP textual encoder and the visual classifiers from the support set. By freezing the CLIP model parameters, only the prompts are optimized during meta-training. In mathematics, the optimization objective can be formulated as:
$\max_{\bm{u}}\mathbb{E}_{p(\tau)}\mathbb{E}_{\bm{x}\sim\mathcal{D}_{\tau}^{Q}}\left[\log p_{\bm{\theta}}(y|\bm{x},\bm{u})\right],$ (48)
where $\bm{u}$ denotes the learnable multi-modal prompts. $\bm{\theta}=(\bm{\theta}_{t},\bm{\theta}_{i})$ contains the parameters of the CLIP textual and visual encoders, which are frozen during training. The prediction of each query image $\bm{x}$ from task $\tau$ is calculated by Eq. (45). Loglikelihood maximization is implemented by minimizing the classification cross-entropy loss.
Many-shot classification.
The many-shot prompt learning setting refers to the 16-shot classification task proposed in work 96, 97. Based on the MaPLe backbone 44, we again freeze the CLIP model parameters and tune multi-modal prompts. The prompts are optimized on a selected ImageNet subset, with 16 samples from each category.
Note that in the many-shot setting, we do not have pre-defined N-way K-shot tasks, either the splits of support and query sets in each task. Therefore, we replace the “tasks” in the meta-learning setting with training samples. Model predictive task sampling is then achieved through data sampling. In mathematics, the objective can be formulated as:
$\max_{\bm{u}}\mathbb{E}_{\bm{x}\sim\mathcal{D}}\left[\log p_{\bm{\theta}}(y|\bm{x},\bm{u})\right],$ (49)
where $\bm{u}$ and $\bm{\theta}$ denote the prompts and frozen CLIP parameters as in the few-shot prompt learning setting, respectively. $\bm{x}$ are training samples from the entire training set $\mathcal{D}$. The prediction of each image $p_{\bm{\theta}}(y|\bm{x},\bm{u})$ is calculated by $\frac{\exp(-d(f_{\bm{\theta}_{i}}(\bm{x},\bm{u}),\bm{t}_{k}))}{\sum_{k^{\prime}}\exp(-d(f_{\bm{\theta}_{i}}(\bm{x},\bm{u}),\bm{t}_{k^{\prime}}))}$ with the textual classifiers $\bm{t}$, obtained similarly to Eq. (43).
Appendix G Experimental Setups & Implementation Details
Practical Learning Efficiency and Robustness.
Widely recognized in reinforcement learning is the high sample complexity in policy evaluation, which demands massive interactions with environments, while policy optimization over the MDP distribution makes this even more severe. In N-way K-shot image classification, we can create K-shot classification task from an arbitrary combination of $N$ classes; then the task space complexity $\mathcal{O}\Big{(}C_{M}^{N}\Big{)}$ grows with the number of categories $M$ in image datasets. Meanwhile, challenges arise when gradient updates of foundation models consume substantial computational power and memory with a large batch size. Similar circumstance also occurs in robust finetuning foundation models. We refer the reader to Supplementary Notes Section F/G for all details of experimental setups and implementations.
Table 4: A Summary of the Considered Benchmarks. Here, we list the primary expensive part in task episodic learning for each scenario together with backbone methods. Also note that N-way K-shot image classification and PEFT requires implicit task identifiers while others can directly access explicit task identifiers as the lower dimensional task representation.
      Benchmarks       Adaptation       Backbone       Expensive Part
      K-shot sinusoid regression       few-shot       MAML       computations
      N-way K-shot image classification       few-shot       MaPLe       computation/memory
      Meta-RL       few-shot       MAML       interactions
      domain randomization       zero-shot       TD3       interactions
      PEFT       many-shot       MaPLe       computation/memory
Neural architecture of the risk learner.
As mentioned in the main paper, the risk learner is in an encoder-decoder structure. For generality sake, we keep the neural architecture same for all benchmarks, including regression, classification and reinforcement learning. The encoder includes an embedding network with 4 hidden layers of size 128 (for Image Classification and Prompt-Tuning) or 10 (for Sinusoid Regression, Meta-RL, and Robotic DR) with the Rectified Linear Unit (ReLU) nonlinear activation units to encode $\{[\bm{\tau}_{t,i},\ell_{t,i}]\}$ batch for mean pooling and then maps to $[\bm{\mu},\bm{\Sigma}]$ with an output layer. The decoder is a network with 3 hidden layers with nonlinear activation units to map $[\bm{z},\bm{\tau}]$ to $\ell\in\mathbb{R}$. For further details, please refer to our code.
Visualized Results during Training Phases.
Note that the active selection and the random sampler with different batches affect the reflection of the machine learner’s performance. Hence, learning curves in sinusoid regression (Fig. 3.a), Meta-RL (Fig. 5.a-b), and DR (Fig. 6.a-b) are actually evaluated in a uniformly sampled validation task dataset for fair comparison. Details on these validation task dataset are attached in the opensourced code.
G.1 Sinusoid Regression
Task setup.
For sinusoid regression, we retain the setup in MAML28, where the few-shot machine learner tries to complete a wave function with the support dataset. In specific, sampling the amplitude $a$ and the phase $b$ configures the wave function, and $10$ data points are uniformly sampled from the interval $[-5.0,5.0]$ coupled with $y=a\sin(x-b)$ to obtain the support dataset. This formulates the $10-\texttt{shot}$ sinusoid regression task.
Meta training process and neural architectures.
The machine learner is a neural network with 2 hidden layers of size $40$ with two nonlinear activation units. The task batch for ERM and GDRM is $16$, and that for DRM is $32$ as default. The temperature parameter in GDRM is $\eta=0.001$. The learning rates for the inner loop and the outer loop are $0.001$. The following is about extra optimization details or setups in MPTS. The task identifier’s dimension is $2$ with the latent variable is $\bm{z}\in\mathbb{R}^{x}$. The batch size of the identifier in training is $32$, the Lagrange multiplier is set as $1$, and we use the Adam optimizer with the learning rate $43e-$ 4 to update the risk learner for 20000 step. In sinusoid regression and Meta-RL, we use the standard repository in MAML 28.
G.2 N-way K-shot Image Classification
Task Setup.
This is a commonly seen benchmark in few-shot learning. It learns a model that can classify images from $N$ distinct classes with support of $K$ labeled examples for each class. The support dataset as reference is in the form $\mathcal{D}_{\tau}^{S}=\{\{[\bm{x}_{i,k},y_{i,k}=i]\}_{k=1}^{K}\}_{i=1}^{N}$. And the query dataset corresponds to the image information for the model to classify. Hence, for a large image dataset with $M$ classes, the complexity of the task space is $\mathcal{O}\Big{(}C_{M}^{N}\Big{)}$. Here, we include ImageNet-CG 47, ImageNet-CI 47, ImageNet-CS 47, ImageNet-A 48, ImageNet-S 49 and ImageNet-R 50 as the dataset in evaluation.
Meta training process and neural architectures.
The machine learner utilizes a prompt learning backbone following MaPLe, with the frozen CLIP model. The task batch for ERM and GDRM is 4, and that for DRM is 8 as default. The temperature parameter in GDRM is 0.001. The learning rate for the outer loop is 0.01. The learning rate for the inner loop follows that in MaPLe 44. The following is about extra optimization details or setups in MPTS. The task identifier is generated by the frozen CLIP text encoder using the input class names, with a dimensionality of 512. The batch size of the identifier in training is 8, the Lagrange multiplier is set as $\beta$, and we use the Adam optimizer with the learning rate 0.01 to update the risk learner.
G.3 Meta-RL
Task Setup.
We construct MDP distributions based on Mujoco physics engines 98. These include HalfCheetahVel, HalfCheetahMassVel, Walker2dVel, Walker2dMassVel, and ReacherPos. The HalfCheetahVel and Walker2dVel tasks involve training the cheetah or walker robot to achieve a target velocity. These tasks define the reward function as the negative absolute difference between the robot’s current velocity and the target velocity, supplemented by a control penalty and an alive bonus to facilitate the learning process. The goals and rewards of HalfCheetahMassVel and Walker2dMassVel are the same as those of the corresponding velocity-related tasks, with the additional identifier of varying mass for the cheetah or walker robot. The ReacherPos task tries to move a two-jointed robot arm’s end effector close to a target position, and its reward function is defined as the negative $L$-1 distance between the robot arm’s position and the target position, supplemented by a control cost to ensure robustness.
Meta training process and neural architectures.
The machine learner is a neural network with 2 hidden layers of size 64 with the Rectified Linear Unit (ReLU) nonlinear activation units. The task batch for ERM and GDRM is 20, and that for DRM is 40 as default. The temperature parameter in GDRM is 0.001. The learning rates for the inner loop and the outer loop are 0.1. The following is about extra optimization details or setups in MPTS. The task identifier is encoded into the latent variable $\bm{z}\in\mathbb{R}^{x}$. The batch size of the identifier in training is 30, the Lagrange multiplier is set as $\beta=0.0001$, and we use the Adam optimizer with the learning rate 0.005 to update the risk learner.
G.4 Robotic DR
Task Setup.
We conduct experiments on LunarLander-v2 and ErgoReacher-v0 environments 9. LunarLander is a 2 degrees of freedom (DoF) environment in which the agent has to land a spacecraft on a designated landing pad without crashing, implemented using Box2D 99. The reward function of LunarLander awards positive rewards for successful landings, negative rewards for crashes, and additional penalties for fuel consumption and deviation from the landing pad, encouraging efficient and controlled landings. ErgoReacher is a 4 DoF arm environment from Golemo et al. 100 in which the arm has to touch a goal with its end effector, implemented in the Bullet Physics Engine 101. The reward function of ErgoReacher includes the negative distance between the end effector’s position and the target, along with other control costs to promote efficient and safe movements. In LunarLander, we randomize the engine strength, while in ErgoReacher, we randomize the joint damping and maximum torque for each of the 4 joints, resulting in a total of 8 parameters. The detailed ranges of the randomized parameters for each environment are provided in Table 3.
DR training process and neural architectures.
The machine learner is a neural network with 2 hidden layers of size 10 with the Rectified Linear Unit (ReLU) nonlinear activation units. The task batch for ERM, and GDRM is 10, and that for DRM is 20 as default. The temperature parameter in GDRM is 0.01. The learning rates for actor and critic are 3e-4. The following is about extra optimization details or setups in MPTS. The task identifier is encoded into the latent variable $\bm{z}$. The batch size of the identifier in training is 25 for LunarLander and 250 for ErgoReacher. The Lagrange multiplier is set as $\beta=1.0$, and we use the Adam optimizer with the learning rate 0.005 to update the risk learner.
G.5 Prompt-Tuning Multimodal Foundation Models as Many-shot Learning
Task Setup of Prompt-tuning.
We refer the reader to MaPLe’s implementation in https://github.com/muzairkhattak/multimodal-prompt-learning. For all baselines, we retain the MaPLe’s task construction in prompt-tuning.
Prompt-tuning process and neural architectures.
The machine learner follows the prompt learning method MaPLe 44 based on the frozen CLIP model (ViT/B-16). The task batch for ERM and GDRM is 4, and that for DRM is 8 as default. The temperature parameter in GDRM is 0.001. The learning rate for the outer loop is 0.005. The learning rate for the inner loop follows that in MaPLe 44. The following is about extra optimization details or setups in MPTS. As for the neural architecture of the risk learner, the encoder is a neural network with 5 hidden layers with 4 ReLU nonlinear activation units, and the decoder is a neural network with 4 hidden layers with 3 nonlinear activation units. The task identifier’s dimension is 512 with the latent embedding from CLIP encoders. The batch size of the identifier in training is 8, the Lagrange multiplier is set as $\beta$, and we use the Adam optimizer with the learning rate 0.005 to update the risk learner for 8000 steps. During prompt-tuning, we set the task batch for ERM and DRM as $\mathcal{B}=4$ (For implementation simplicity, the data loader samples $8$ tasks and then randomly keeps half without ranking to optimize). The task batch for DRM is $\mathcal{B}=8$ before the filtering operation; DRM filters half to optimization. Similarly, that for MPTS is $\mathcal{B}=8$ and only $4$ tasks are screened to optimize.
Table 5: Testing Classification Results after 5-way 1-shot Meta-Training on Various Datasets. We report testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with the meta-trained machine learner on ID and OOD datasets. This table complements the radar part of Fig. 4a-d. Best results are in bold, and MPTS’s performance gains over ERM are marked in blue.
      Dataset       Metrics       ERM       DRM       GDRM       MPTS (Ours)
      ImageNet-CG 47       $\text{CVaR}_{0.9}$       77.02       77.76       77.61       78.04       +1.02
      $\text{CVaR}_{0.7}$       82.00       82.47       82.62       82.87       +0.87
      $\text{CVaR}_{0.5}$       84.77       85.03       85.16       85.45       +0.68
      Average       89.04       89.46       89.51       89.87       +0.83
      ImageNet-CI 47       $\text{CVaR}_{0.9}$       80.24       80.47       80.15       80.97       +0.73
      $\text{CVaR}_{0.7}$       84.77       85.17       84.78       85.66       +0.98
      $\text{CVaR}_{0.5}$       87.03       87.52       87.09       87.78       +0.75
      Average       91.12       91.46       91.15       91.60       +0.48
      ImageNet-CS 47       $\text{CVaR}_{0.9}$       76.63       77.71       76.18       78.4       +1.77
      $\text{CVaR}_{0.7}$       81.58       82.53       81.40       83.27       +1.69
      $\text{CVaR}_{0.5}$       84.40       85.08       84.23       85.83       +1.43
      Average       89.24       89.87       89.13       90.26       +1.02
      ImageNet-A 48       $\text{CVaR}_{0.9}$       76.69       77.39       77.40       77.95       +1.26
      $\text{CVaR}_{0.7}$       81.90       82.57       82.50       83.41       +1.51
      $\text{CVaR}_{0.5}$       84.58       85.43       85.33       86.30       +1.72
      Average       89.25       90.26       90.21       91.06       +1.81
      ImageNet-S 49       $\text{CVaR}_{0.9}$       82.63       83.67       83.11       84.82       +2.21
      $\text{CVaR}_{0.7}$       87.26       88.38       87.26       89.27       +2.01
      $\text{CVaR}_{0.5}$       89.56       90.02       90.08       91.54       +1.98
      Average       93.53       94.27       94.12       94.78       +1.25
      ImageNet-R 50       $\text{CVaR}_{0.9}$       88.31       88.84       87.71       89.79       +1.48
      $\text{CVaR}_{0.7}$       91.46       92.03       91.18       93.16       +1.7
      $\text{CVaR}_{0.5}$       93.36       93.80       93.21       94.65       +1.29
      Average       96.05       96.33       95.98       96.86       +0.81
Table 6: Testing Classification Results after Prompt-Tuning on ImageNet. We report testing $\text{CVaR}_{0.9}$, $\text{CVaR}_{0.7}$, $\text{CVaR}_{0.5}$ and average accuracies with the prompt-tuned machine learner on ID and OOD datasets. Evaluation on OOD datasets corresponds to the domain generalization setting. This table complements the radar part of Fig. 7a-d. Best results are in bold, and MPTS’s performance gains over ERM are marked in blue.
     Dataset      Metrics      ERM      DRM      GDRM      MPTS (Ours)
     ImageNet 51 (ID)      $\text{CVaR}_{0.9}$      31.68      32.46      31.38      32.5      +0.82
     $\text{CVaR}_{0.7}$      42.87      44.07      42.97      44.22      +1.35
     $\text{CVaR}_{0.5}$      51.45      52.59      51.71      52.72      +1.27
     Average      70.8      70.8      71.0      71.20      +0.4
     ImageNet-A 48 (OOD)      $\text{CVaR}_{0.9}$      15.33      15.46      15.6      18.44      +3.11
     $\text{CVaR}_{0.7}$      22.8      23.02      23.13      24.06      +1.22
     $\text{CVaR}_{0.5}$      30.08      29.54      30.26      31.25      +1.17
     Average      49.8      48.4      49.5      51.10      +1.3
     ImageNet-R 50 (OOD)      $\text{CVaR}_{0.9}$      26      28.06      25.9      28.19      +2.19
     $\text{CVaR}_{0.7}$      43.58      45.23      43.91      45.49      +1.91
     $\text{CVaR}_{0.5}$      56.67      58.21      57.38      58.77      +2.1
     Average      76.9      77.4      77.3      77.63      +0.73
     ImageNet-S 49 (OOD)      $\text{CVaR}_{0.9}$      12.24      13.16      12.22      13.63      +1.39
     $\text{CVaR}_{0.7}$      20.02      21.08      20.43      21.46      +1.44
     $\text{CVaR}_{0.5}$      26.69      27.44      27.4      27.97      +1.28
     Average      48.9      48.8      48.9      49.63      +0.73
     ImageNet-V 52 (OOD)      $\text{CVaR}_{0.9}$      24.6      25.5      24.2      25.90      +1.3
     $\text{CVaR}_{0.7}$      34.7      35.73      35.07      35.80      +1.1
     $\text{CVaR}_{0.5}$      43.45      44.14      43.53      44.56      +1.11
     Average      64.0      63.8      64.1      64.53      +0.53
Appendix H Computational Tools & Platforms & Data Availability
In this research project, we use the Pytorch as the package to implement all methods to run all deep learning epxeriments.
Appendix I Competing Interests & Author Contributions
The author list is Qi Cheems Wang (Q.C.W.), Zehao Xiao (Z.X.), Yixiu Mao (Y.M.), Yun Qu (Y.Q.), Jiayi Shen (J.S.), Yiqin Lv (Y.L.), and Xiangyang Ji (X.J.). The authors declare no competing interests in this work. X.J. launched and sponsored this research on the reliable and efficient adaptation learning project. Z.X. and J.S. from the University of Amsterdam attended this project, and this work was done during their remote visiting the Tsinghua University Intelligent Decision-Making Lab from April 2024 to August 2024. J.S. is now working at Facebook AI Research. The authors confirm their contributions to this work as follows:
Under the supervision of professor X.J., Q.C.W. conceptualized the idea of MPTS, designed the computational framework, formulated the mathematical part, and wrote the draft. Z.X. and J.S. implemented MPTS in sinusoid regression, prompt-based few-shot image classification, and PEFT image classification, collected experimental results to visualize, and added implementation details in Supplementary Material. Y.M., Y.Q., and Y.L. implemented MPTS in sinusoid regression, Meta-RL and robotic DR, collected experimental results to visualize, and added implementation details in Supplementary Material. X.J. supervised the progress of MPTS, organized technical discussions, reviewed and revised the original draft. All authors have read the manuscript and approved the public version.
First Author Biography:
Qi Wang received Ph.D. degree under supervision of Professor Max Welling and Associate Professor Herke van Hoof in 2022. He is now under supervision of Prof. Xiangyang Ji and works as a research assistant at Tsinghua University. His research focus is on generative modeling and intelligent decision-making. He has published several papers on top-tier conferences such as ICML/NeurIPS/ICLR and was awarded 2023 China Multi-Agent System Outstanding Doctoral Thesis Award.
Correspondence Author Biography:
Xiangyang Ji received the B.S. degree in materials science and the M.S. degree in computer science from the Harbin Institute of Technology, Harbin, China, in 1999 and 2001, respectively, and the Ph.D. degree in computer science from the Institute of Computing Technology, Chinese Academy of Sciences, in 2008. He joined Tsinghua University, Beijing, in 2008, where he is currently a Professor with the Department of Automation, School of Information Science and Technology. He has authored more than 100 refereed conference and journal papers. His current research interests include signal processing, computer vision, computational photography, and intelligent decision-making.